### v1

# Distribution Learning Meets Graph Structure Sampling

Abstract
This work establishes a novel link between the problem of PAC-learning high-dimensional graphical models and the task of (efficient) counting and sampling of graph structures, using an online learning framework. The problem of efficiently counting and sampling graphical structures, such as spanning trees and acyclic orientations, has been a vibrant area of research in algorithms. We show that this rich algorithmic foundation can be leveraged to develop new algorithms for learning high-dimensional graphical models.
We observe that if we apply the exponentially weighted average (EWA) or randomized weighted majority (RWM) forecasters on a sequence of samples from a distribution $P^{*}$ using the log loss function, the average regret incurred by the forecaster’s predictions can be used to bound the expected KL divergence between $P^{*}$ and the predictions. Known regret bounds for EWA and RWM then yield new sample complexity bounds for learning Bayes nets. Moreover, these algorithms can be made computationally efficient for several interesting classes of Bayes nets.
  1. (i)
    For the class of tree-structured distributions over $[k]^{n}$, we design an algorithm that, for any $\varepsilon>0$, uses $\widetilde{O}(nk^{2}/\varepsilon)$ samples from some distribution $P^{*}$ and returns $\widehat{P}$, an efficiently samplable mixture of trees, whose KL-divergence from $P^{*}$ is at most 3 times the optimal KL-divergence between $P^{*}$ and any tree-structured distribution plus $\varepsilon$, with probability at least $2/3$. This guarantee is not known to be achievable with the well-known Chow-Liu algorithm.
  2. (ii)
    For any constant $d>0$, given an undirected chordal graph $G$, we design an algorithm that uses $\widetilde{O}(n^{3}k^{d+1}/\varepsilon^{2})$ samples from some distribution $P^{*}$ over $[k]^{n}$, runs in polynomial time, and returns a Bayes net $\widehat{P}$ with skeleton $G$ and indegree $d$, whose KL-divergence from $P^{*}$ is at most $\varepsilon$ more than the optimal KL-divergence between $P^{*}$ and any Bayes net with skeleton $G$ and indegree $d$, with probability at least $2/3$. The only previous result in this spirit was for $G$ being a tree and in the realizable setting.
1 Introduction
High-dimensional distributions are foundational to contemporary machine learning, with applications that span across various domains including gene regulation networks [40, 24, 25, 39], protein signaling networks [29, 69, 74], brain connectivity networks [49, 77], and psychiatric symptom networks [12, 63, 79] etc. In particular, probabilistic graphical models, including Bayesian networks, Ising models, and Gaussian graphical models, are extensively utilized to model a wide range of data generation processes in practice (see [58, 78, 52] and the references therein). These models serve as powerful tools for understanding complex relationships and dependencies within high-dimensional data sets, making them indispensable in both theoretical and applied machine learning. A graphical model is a succinct representation of a high-dimensional distribution over an (exponentially) large sample space such as $\mathbb{R}^{n}$ or $\{0,1\}^{n}$. Due to the limited dependence between component random variables encoded by a dependence graph, these models are particularly powerful for describing the joint probability distribution over a large set of variables in a succinct and interpretable way. Learning distributions represented by graphical models is a central problem with both theoretical and practical significance.
A Bayesian network (Bayes net) with $n$ variables and alphabet size $k$ is a probability distribution over the sample space $[k]^{n}$ defined with respect to a directed acyclic dependency graph (DAG) $G$ over the vertex set $[n]$. Each node of $G$ defines a component random variable over the sample space $[k]$. Every component $i$ is independent of all the components that are not descendants of $i$ conditioned on the components that are the parents of $i$. Using Bayes rule, it can be shown that such a distribution then factorizes as a product of $n$ conditional probability values. Thus if $G$ has indegree at most $d$, such a distribution can be encoded by at most $nk^{d+1}$ parameters instead of $k^{n}$.
Learning Bayesian network distributions from samples generally involves two steps: structure learning, where the underlying dependency graph is identified, and parameter learning, which involves determining the conditional probability table when the structure is known. Structure learning methods fall into two categories: constraint-based and score-based. Constraint-based approaches begin with a full directed acyclic graph (DAG) and eliminate edges by testing conditional independences, requiring significant conditional dependencies for accurate graph recovery. Score-based methods, on the other hand, assign scores to each DAG and formulate structure recovery as an optimization problem, often tackled with local heuristics like greedy hill climbing due to the complexity of the optimization.
The current work reprises the score-based approach towards Bayes net learning, but instead of optimizing the score directly, we use the framework of online learning to reduce the problem to sampling from a family of high-dimensional structures. In the setup of online learning, the goal is to design a forecaster that observes a sequence of examples $x^{(1)},x^{(2)},\dots,x^{(T)}$, and at each time $t$, outputs a prediction $\widehat{p}_{t}$ based only on $x^{(1)},\dots,x^{(t-1)}$. After predicting $\widehat{p}_{t}$, it observes $x^{(t)}$, and it incurs a loss $\ell(\widehat{p}_{t},x^{(t)})$ for a loss function $\ell$. The cumulative loss of the forecaster is benchmarked against that of a fixed and known set of experts. The goal of the algorithm is to minimize the regret, defined as the difference between the cumulative loss over all rounds and the loss it would incur if it were to follow the best expert in hindsight. Online learning is a mature field with numerous applications to theoretical computer science, such as in game theory, approximation algorithms, and complexity theory (see [42, 43, 9, 30, 66, 4, 55] and the references therein).
In the context of distribution learning, we set the set of experts as all possible candidate Bayesian networks (up to a sufficient discretization). The inputs will be random samples from the input distribution, and we set the loss function to be the negative log-likelihood $\ell(\widehat{p},x)=-\log\widehat{p}(x)$. The primary obstacle in applying the online approach to distribution learning lies in ensuring computational efficiency. All the standard forecasting algorithms have running time at least linear in the number of experts; in our case, the experts will consist of all the discretization of all candidate Bayes nets, which is exponentially many. A key insight of our work is the close relationship between this challenge and the task of efficient counting and sampling of DAGs. Recognizing this link allows us to transfer techniques and algorithms from the counting and sampling literature to the realm of distribution learning, leading to significant new results in learning Bayes net distributions.
1.1 Our Results
We first set up the framework of PAC-learning [76] for distributions; formal definitions appear in Section 2. We use KL-divergence as the notion of similarity between probability distributions, and we will throughout work with distributions on $[k]^{n}=\{1,\dots,k\}^{n}$. Let $\mathcal{C}$ be a class of such distributions; in our applications, $\mathcal{C}$ will correspond to some family of Bayes nets.
For $\varepsilon>0,A\geq 1$ and two distributions $P$ and $\widehat{P}$ over $[k]^{n}$, we say $\widehat{P}$ is an $(\varepsilon,A)$-approximation for $P$ with respect to $\mathcal{C}$ if $\mathsf{D_{KL}}(P\|\widehat{P})\leq A\cdot\min_{Q\in\mathcal{C}}\mathsf{D_{KL}}(P\|Q)+\varepsilon$. When $A=1$, we simply say $\widehat{P}$ is an $\varepsilon$-approximation of $P$. An algorithm is said to be an agnostic PAC-learner for $\mathcal{C}$ if for any $\varepsilon,\delta>0$ and access to i.i.d. samples from an input distribution $P^{*}$, it outputs a distribution $\widehat{P}$ which is an $\varepsilon$-approximation for $P^{*}$ with probability at least $1-\delta$. If $\widehat{P}$ is not necessarily in $\mathcal{C}$, the algorithm is called improper; otherwise, it is called proper. Also, the realizable setting corresponds to when the input $P^{*}$ is guaranteed to be in $\mathcal{C}$.
General Framework for Learning Bayes Nets
It is well-known (e.g., [8]) that given a DAG $G$, the minimum KL divergence between $P^{*}$ and a Bayes net over $G$ can be written as $J_{P^{*}}-\sum_{i\in V(G)}I(X_{i};X_{\mathsf{Pa}_{G}(i)})$, where $X\sim P^{*}$ and $J_{P^{*}}$ is a constant independent of $G$. Hence, if $\mathcal{C}$ is the class of Bayes nets over DAGs of in-degree $d$, a natural strategy for designing agnostic learning for $\mathcal{C}$ is to approximate the mutual information between any node and any set of $d$ other nodes up to a suitable additive error. Next, one maximizes over all possible DAGs with in-degree $d$, the sum of mutual informations between a node and its $d$ parents. Iterating over all possible DAG structures would lead to an algorithm with sample complexity $\widetilde{O}(n^{2}k^{d+1}\varepsilon^{-2})$ and exponential time complexity.
Our proposed approach immediately gives an improper learning guarantee with a quadratically better sample complexity in the realizable setting (but again with exponential time complexity).
Theorem 1.1.
Let $\mathcal{C}$ be the family of distributions over $[k]^{n}$ that can be defined as Bayes nets over DAGs with $n$ nodes and maximum in-degree $d$. There exists a PAC-learner for $\mathcal{C}$ in the realizable setting using ${O}(nk^{d+1}\varepsilon^{-1}\log(n/\varepsilon))$ samples that returns a mixture of Bayes nets from $\mathcal{C}$ which is an $\varepsilon$-approximation of the input distribution with probability at least $2/3$.
The sample complexity in 1.1 is nearly optimal [^1] Note that for learning with respect to total variation distance, the corresponding sample complexity of $\widetilde{O}(nk^{d+1}\varepsilon^{-2})$ is known [5], but the proof uses the tournament method which does not apply to KL divergence as it is not a metric. for constant error probability. The same approach also yields polynomial sample complexity for agnostic learning in the non-realizable setting with exponentially small error probability and for proper learning with polynomially small error probability. While these bounds are no longer optimal, we will use them later to design the first polynomial time algorithms for learning Bayes nets over large classes of DAGs.
Learning Tree-structured Distributions
By tree-structured distributions (or simply, trees when the meaning is clear), we mean Bayes nets whose underlying DAG has in-degree 1. They can be equivalently defined as undirected Markov models over trees. Prior works have established tight sample complexity bounds for learning trees. It is known [38, 8] that for proper PAC-learning in the realizable setting, $\widetilde{\Theta}(n\varepsilon^{-1})$ samples are necessary, that for agnostic proper PAC-learning $\widetilde{\Theta}(n^{2}\varepsilon^{-2})$ samples are necessary, and that both bounds are nearly achieved by the classic Chow-Liu algorithm [22].
In our proposed framework, we surprisingly get a learning algorithm for $(\varepsilon,3)$-approximating trees using $\widetilde{\Theta}(nk^{2}\varepsilon^{-1})$ samples in the agnostic (non-realizable) setting. Note that this is quadratically better than the aforesaid optimal bound for agnostic proper learning of trees.
Theorem 1.2.
Let $\mathcal{C}$ be the family of tree-structured distributions over $[k]^{n}$. There exists an algorithm that for any $\varepsilon>0$, given sample access to a distribution $P^{*}$, returns an $(\varepsilon,3)$-approximation $\widehat{P}$ of $P^{*}$ with probability at least $2/3$ and uses $m=\widetilde{O}(nk^{2}\varepsilon^{-1})$ samples and $\mathrm{poly}(m)$ running time. Moreover, there is a polynomial time algorithm to generate samples from $\widehat{P}$.
A curious corollary is that in the realizable setting, we get a PAC-learning algorithm for learning trees with sample complexity $\widetilde{O}(nk^{2}\varepsilon^{-1})$ that is very different from Chow-Liu. In fact, the current best upper bound on the sample complexity of Chow-Liu from [8] is $\widetilde{O}(nk^{3}\varepsilon^{-1})$, which is worse by a factor of $k$. The output distribution $\widehat{P}$ is a mixture of exponentially many trees but can nevertheless be sampled in polynomial time by using the matrix-tree theorem, as we explain later. Learning with mixtures of trees has been studied before [60, 53, 1, 73] but in other contexts.
We also note that going beyond trees, the same approach allows us to develop polynomial sample and time algorithms for learning Bayes nets on an unknown DAG whose moralization is promised to have constant vertex cover size. Here, instead of sampling using the matrix-tree theorem, we can utilize a recent result of Harviainen and Koivisto [47] to sample such DAGs. Details appear in Appendix B.
Learning with Known Chordal Skeleton
The skeleton of a DAG refers to its underlying undirected graph. A well-investigated class of Bayes nets that go beyond trees are polytrees, i.e., Bayes nets whose DAGs have tree skeletons. For example, a Bayes net on an inward directed star is a polytree but not a tree. Polytrees are especially interesting because they admit fast exact inference [62]. Dasgupta [28] looked at the problem of learning polytree distributions in terms of the negative log-likelihood cost. He showed that it is NP-hard to get a 2-polytree (Bayes nets of indegree 2 whose skeleton is acyclic) whose cost is at most $c$ times that of the optimal 2-polytree for some constant $c>1$, even if we have oracle access to the true entropies (equivalently, infinite samples). Our reverse-KL cost differs from Dasgupta’s cost by an additive term that corresponds to the entropy of the true unknown distribution [8, Lemma 3.3]. Moreover, our distribution learning algorithms achieve an additive approximation in the reverse-KL cost.
More generally, we consider Bayes nets having a chordal skeleton with bounded indegree; polytrees with bounded indegree are a special case.
Theorem 1.3.
Let $G$ be an undirected chordal graph, and suppose $k$ and $d$ are fixed constants. Let $\mathcal{L}^{G}_{d}$ be the family of distributions over $[k]^{n}$ that can be defined as Bayes nets over DAGs having skeleton $G$ and with indegree at most $d$. There exist (i) an agnostic PAC-learner for $\mathcal{L}^{G}_{d}$ using $\widetilde{O}(n^{4}k^{2d+2}\varepsilon^{-4}\log\delta^{-1})$ samples that is improper and returns an efficiently-samplable mixture of distributions from $\mathcal{L}^{G}_{d}$, and (ii) an agnostic PAC-learner for $\mathcal{L}^{G}_{d}$ using $\widetilde{O}(n^{3}k^{d+1}\varepsilon^{-2}\log(n/\varepsilon\delta))$ samples and polynomial running time that is proper and returns a distribution from $\mathcal{L}^{G}_{d}$.
The closest related result is that of [26] who designed a PAC-learner in the realizable setting for polytrees with optimal sample complexity $\widetilde{O}(nk^{d+1}\varepsilon^{-1}\log\delta^{-1})$. However, their analysis crucially uses the realizability assumption, and it was left as an open question in that work to find an efficient agnostic learner for polytrees. The sample complexity bounds in 1.3 follow from the general bounds for learning Bayes nets; the crux of the theorem is the computational efficiency. As we explain below, this is accomplished using a dynamic programming approach, motivated by prior work on counting the number of acyclic orientations in chordal graphs [70, 11, 21].
1.2 Our Techniques
Online Learning Framework to Learning in reverse KL
Given i.i.d. samples from a distribution $P^{*}$, we are trying to learn it. Roughly, for a random $x\sim P^{*}$, a good approximate Bayes net $P$ should maximize the probability $P(x)$, or equivalently, minimize the expected log-likelihood $\mathbb{E}_{x\sim P^{*}}\left[\log\frac{1}{P(x)}\right]$. Keeping this in mind, we define the loss of any Bayes net $\widehat{P}$ predicted by the algorithm to be $\log\frac{1}{\widehat{P}(x)}$ for a sample $x$.
We follow the online learning framework. Here the algorithm $\mathcal{A}$ observes a set of samples $x^{(1)},\dots,x^{(T)}\sim P^{*}$ over $T$ rounds from an unknown Bayes net $P^{*}$. The goal of the algorithm is to learn a distribution $\widehat{P}$ which is close to $P^{*}$. After observing each sample $x^{(t)}$, $\mathcal{A}$ predicts a Bayes net $\widehat{P}_{t}$ and incurs a loss of $\log\frac{1}{\widehat{P}_{t}(x^{(t)})}$ for this round. However, there is a set of experts $\mathcal{E}$ to help $\mathcal{A}$. For simplicity, we can assume each expert $E\in\mathcal{E}$ is simply one Bayes net among all possible Bayes nets. Had $\mathcal{A}$ stuck to any particular expert $E\in\mathcal{E}$, it would incur a total loss $\sum_{t=1}^{T}\log\frac{1}{E(x^{(t)})}$ over all the $T$ rounds. The algorithm can change the experts in between or do some randomized strategy for choosing the expert among $\mathcal{E}$. Let $\widehat{P}_{t}$ be its prediction after round $t$. The regret is defined to be the difference between the loss of the algorithm and that of the best expert: $\sum_{t=1}^{T}\log\frac{1}{E(x^{(t)})}-\min_{E\in\mathcal{E}}\sum_{t=1}^{T}\log\frac{1}{E(x^{(t)})}$.
We design a set of finitely many Bayesian networks that “covers” all Bayes nets to get our set of experts. This discretization incurs a loss that we can control (see Section 3.2 for the details). Importantly, we relate the regret mentioned above to the expected average of the reverse KL divergence over the rounds: $\mathbb{E}[\frac{1}{T}\mathsf{D_{KL}}(P^{*}||\widehat{P}_{t})]$. Learning in reverse KL divergence $\mathsf{D_{KL}}(P^{*}||\cdot)$ has implications for maximum likelihood and learning in metric distance functions such as the total variation and Hellinger distance. Therefore, learning in the reverse KL divergence has received considerable attention in machine learning. Once we control the average distance, using the convexity of KL, we can show that the mixture distribution $\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}$ is also close to $P^{*}$. Finally, we manage to translate the above bounds from expectation to high probability using the bounded difference inequality due to McDiarmid.
In particular, we use two different online algorithms: the Exponential Weighted Average (EWA) algorithm and the Randomized Weighted Majority (RWM) algorithm. EWA returns us the mixture $\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}$ which improperly learns $P^{*}$ in (reverse) KL. RWM returns a random Bayes net $\widehat{P}$ which properly learns $P^{*}$ in expected KL.
Efficient Learning of Restricted Classes of Bayes Nets
Our learning algorithm for Bayes nets mentioned above is sample-optimal but not time-efficient in general, since the number of experts to be maintained is of exponential size. However, we observe that for special cases of Bayes nets, we can efficiently sample from the experts according to the randomized strategy of the algorithm. As a remark, the idea that the computational barrier of RWM or EWA may be side-stepped by developing efficient sampling schemes which were also used in a recent work on fast equilibrium computation in structured games [9] and partly motivated our work.
To see the simplest example of this idea, suppose $\mathcal{P}=\{P_{1},\dots,P_{N}\}$ is a set of distributions over $[k]$, and let $\mathcal{P}^{\otimes n}=\mathcal{P}\times\mathcal{P}\times\cdots\times\mathcal{P}$ be a set of product distributions over $[k]^{n}$. Each element of $\mathcal{P}^{\otimes n}$ is indexed as $P_{\mathbf{i}}$ for $\mathbf{i}=(i_{1},\dots,i_{n})$, so that $P_{\mathbf{i}}(x)=\prod_{j=1}^{n}P_{i_{j}}(x_{j})$. The size of $\mathcal{P}^{\otimes n}$ is clearly $N^{n}$, so it is infeasible to work with it directly. The RWM algorithm maintains a distribution over $\mathcal{P}^{\otimes n}$, so that the probability that RWM picks $P_{\mathbf{i}}$ for its prediction $\widehat{P}_{t}$ at time $t$ is proportional to $\prod_{s=1}^{t-1}P_{\mathbf{i}}(x^{(s)})^{\eta}$, where $x^{(s)}$ is the observed sample at time $s$ and $\eta>0$ is a parameter. Therefore:
$\Pr_{\textsf{RWM}}[\widehat{P}_{t}=P_{\mathbf{i}}]=\frac{\prod_{s=1}^{t-1}P_{\mathbf{i}}(x^{(s)})^{\eta}}{\sum_{\mathbf{i}^{\prime}}\prod_{s=1}^{t-1}P_{\mathbf{i}^{\prime}}(x^{(s)})^{\eta}}=\prod_{j=1}^{n}\frac{\prod_{s=1}^{t-1}P_{i_{j}}(x^{(s)}_{j})^{\eta}}{\sum_{i^{\prime}_{j}}\prod_{s=1}^{t-1}P_{i^{\prime}_{j}}(x^{(s)}_{j})^{\eta}}.$
The crucial observation is that RWM maintains a product distribution over product distributions, and so we can sample each $P_{i_{j}}$ from $\mathcal{P}$ independently.
When the underlying Bayes net is a tree, i.e. of indegree 1, we show that RWM samples a random rooted spanning arborescence from a weighted complete graph. The probability to output a particular arborescence $A$ is proportional to $\prod_{e\in A}w_{e}$ where each $w_{e}$ is a weight that can be explicitly computed in terms of the observed samples and the parameters of the algorithm. It is well-known that the matrix-tree theorem (more precisely, Tutte’s theorem) for counting weighted arborescences can be used for this purpose, and hence, we obtain an alternative to the Chow-Liu algorithm for approximately learning a tree Bayes net efficiently and sample-optimally.
[…]
For simplicity, let us assume Boolean alphabet, i.e. $k=2$. As discussed before, due to the factorization property of Bayes nets, given a fixed orientation $T$ of the path skeleton, iterating over all possible (discretized) conditional probabilities for each of the $n$ variables is as good as iterating over $T$-structured Bayes nets. For path graphs, the indegree is either 0, 1, or 2. With a slight abuse of notation, let $\mathcal{B}_{0}$ be the set of all discretized distributions $P(x)$ for a binary variable $x$; let $\mathcal{B}_{1}$ be the set of all discretized distributions $P(x\mid y)$ for binary variables $x,y$; let $\mathcal{B}_{2}$ be the set of all discretized distributions $P(x\mid y,z)$ for binary variables $x,y,z$.
We now show how to compute $Z$ by induction on the set of vertices of the path. Suppose $Z_{j}$ is the normalization constant obtained by only restricting to the first $j+1$ nodes in the path. That is, if $\mathcal{P}_{j}$ is the class of all discretized Bayes nets on $j+1$ variables with a path skeleton, then $Z_{j}=\sum_{P\in\mathcal{P}_{j}}e^{-\eta\ell_{P}}$, where $\ell_{P}$ only computes the loss based on the first $j+1$ variables. For the induction, we maintain more refined information for each $j$. Let $\mathcal{P}_{j,\leftarrow}$ and $\mathcal{P}_{j,\rightarrow}$ be the class of all discretized Bayes nets on $j+1$ variables with a path skeleton and the last edge pointing left and right, respectively. Correspondingly, define $Z_{j,\leftarrow}$ and $Z_{j,\rightarrow}$; clearly, $Z_{j}=Z_{j,\leftarrow}+Z_{j,\rightarrow}$. As the inductive hypothesis, assume that $Z_{j,\leftarrow}$ and $Z_{j,\rightarrow}$ are already computed. We now need to compute $Z_{j+1,\leftarrow}$ and $Z_{j+1,\rightarrow}$.
If the $(j+1)$-th edge orients rightward, then the parents of nodes $11,\dots,j+$ 1 do not change, while the new node $j+2$ has parent $j+1$. We can accommodate this new edge by simply adding the negative log of the conditional probabilities due to this new edge to the loss restricted to the first $j+1$ variables. Computing
$\Delta=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{1}}\log P^{(B)}(x^{(t)}_{j+2}\mid x^{(t)}_{j+1})^{-1},$
we will have
$Z_{j+1,\rightarrow}=(Z_{j,\leftarrow}+Z_{j,\rightarrow})e^{-\eta\Delta}.$
If the $(j+1)$-th edge orients leftward, the adjustment is slightly trickier as node $j+1$ will get a new parent $j+2$, while the new node $j+2$ has no parent. In that case, we need to first subtract out the previous sum of negative log conditional probabilities at $j+1$. Let us define:
$\Delta_{1}=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{1}}\log P^{(B)}(x^{(t)}_{j+1}\mid x^{(t)}_{j})^{-1},$ $\Delta_{2}=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{0}}\log P^{(B)}(x^{(t)}_{j+1})^{-1},$
$\Delta_{3}=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{2}}\log P^{(B)}(x^{(t)}_{j+1}\mid x^{(t)}_{j},x^{(t)}_{j+2})^{-1},$ $\Delta_{4}=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{1}}\log P^{(B)}(x^{(t)}_{j+1}\mid x^{(t)}_{j+2})^{-1},$
$\Delta_{5}=\sum_{t=1}^{T}\sum_{B\in\mathcal{B}_{0}}\log P^{(B)}(x^{(t)}_{j+2})^{-1}.$
If node $j$ is not a parent of node $j+1$, then node $j+1$ contributed $\Delta_{2}$ loss to $Z_{j,\leftarrow}$ while now it contributes $\Delta_{4}$ loss to $Z_{j+1,\leftarrow}$. Otherwise, it contributed $\Delta_{1}$ loss to $Z_{j,\rightarrow}$ while now it contributes $\Delta_{3}$ loss to $Z_{j+1,\leftarrow}$. The new node $j+2$ contributes $\Delta_{5}$ loss to $Z_{j+1,\leftarrow}$ independent of what happens to the other variables. Summarizing:
$Z_{j+1,\leftarrow}=Z_{j,\leftarrow}\cdot e^{-\eta(\Delta_{4}-\Delta_{2}+\Delta_{5})}+Z_{j,\rightarrow}\cdot e^{-\eta(\Delta_{3}-\Delta_{1}+\Delta_{5})}.$
It is easy to see that these updates can be performed efficiently using an appropriate dynamic programming table. Once we have computed the total sum $Z=Z_{n,\leftarrow}+Z_{n,\rightarrow}$, sampling a structure according to the sampler’s distribution can simply be done by suitably unrolling the DP table.
{subcaptionblock}
0.24 Refer to caption {subcaptionblock}0.24 Refer to caption {subcaptionblock}0.24 Refer to caption {subcaptionblock}0.24 Refer to caption
Figure 1: Given a rooted polytree skeleton, for each node $v$, and for each fixed orientation of edges incident to $v$, we maintain the total weight of all consistent orientations of the subtree rooted at $v$. Above, the orientations of edges incident to $B$ and $C$ are fixed. This is needed when computing the weight for the subtree rooted at $A$, since in the first two panels, the in-degree of $C$ change from 1 to 2, while in the second two panels, $C$’s in-degree does not change.
The argument described above extends to learning bounded indegree polytrees and bounded indegree chordal graphs. For polytrees, the idea is illustrated in Figure 1. For chordal graphs, the algorithm first builds a clique tree decomposition and uses this structure for dynamic programming. The obvious issue with chordal graphs is that some orientations may lead to cycles, unlike the case for polytrees. However, chordal graphs enjoy certain nice property (see 5.6) that allows us to independently perform weighted counting/sampling of acyclic orientations in each subtree of the clique tree.
1.3 Related Works
Hoffgen [50] gave the first sample complexity bounds for agnostic learning of a Bayes net with known structure from samples in KL divergence. This work also gave an efficient algorithm for special cases such as trees using the classical Chow-Liu algorithm. Subsequently, Dasgupta [27] gave an efficient algorithm for learning an unknown Bayes net (discrete and Gaussian) on a fixed structure. This result was improved to a sample-optimal learning of fixed-structure Bayes nets in [7, 8].
The general problem of distribution learning of Bayes networks with unknown DAG structure has remained elusive so far. It has not been shown to be NP-hard, although some related problems and specific approaches are NP-hard [17, 20, 36, 56]. Many of the early approaches required faithfulness, a condition which permits learning of the Markov equivalence class, e.g. [67, 18, 41]. Finite sample complexity of such algorithms has also been studied, e.g. [44]. Specifically for polytrees, [64, 46] studied recovery of the DAG for polytrees under the infinite sample regime and in the realizable setting, while [26] gave finite sample complexity for this problem. [45] studied the more general problem of learning Bayes nets, and their sufficient conditions simplified in the setting of polytrees.
A notable prior work in the context of the current paper is the work by Abbeel, Koller, and Ng [2], which also explores the improper learning of Bayesian networks with polynomial sample and time complexities. However, our research diverges from theirs in three critical ways: firstly, their study does not offer any proper learning algorithms; secondly, it lacks agnostic learning guarantees; and thirdly, their approach does not achieve optimal sample complexity in the realizable setting. While they do demonstrate a “graceful degradation” as inputs deviate from the hypothesis class, this does not equate to a definitive agnostic learning guarantee as provided in our work. On a positive note, their research does attain polynomial sample and time complexities for learning any Bayesian network with a bounded total degree in the realizable setting. It is worth noting that our results for distributions with chordal skeleton are applicable even when the total degree is unbounded, provided that the indegree remains bounded, a scenario where the findings of Abbeel, Koller, and Ng would not be applicable.
Online Learning of Structured Distributions
To the best of our knowledge, our proposed approach of learning high-dimensional distributions through online learning has not been reported in the literature before. The closest we get is the Sparsitron algorithm by Klivans and Meka [54] which learns an unknown Ising model from samples. However, Sparsitron is typical to Ising models where the conditional distribution at any component follows a logistic regression model which the Sparsitron algorithm learns.
Robust Learning
In the field of distribution learning, it is commonly assumed that all samples are consistently coming from an unknown distribution. However, real-world conditions often challenge this assumption, as samples may become corrupted—either removed or substituted by samples from an alternate distribution. Under such circumstances, the theoretical assurances of traditional algorithms may no longer apply. This discrepancy has spurred interest in developing robust learning algorithms capable of tolerating sample corruption. Recent years have seen notable advancements in this area, including the development of algorithms for robustly learning Bernoulli product distributions [33], and enhancing the robustness of learning Bayes nets [15]. See [59, 31, 32, 6, 48, 57, 32, 34, 23, 19] and the references therein for a sample of current works in this area. These works primarily focus on guarantees with respect to the total variation distance.
Of particular relevance to us is the TV-contamination model. Here, if the distribution to be learnt is $P$, one gets samples from a ‘contaminated’ distribution $Q$ such that $\mathsf{d_{TV}}(P,Q)\leq\eta$. Note that this is a stronger model of contamination than Huber contamination [51], where the noise is restrictive to be additive, meaning that an adversary adds a limited number of noisy points to a set of uncontaminated samples from $P$.
One can interpret our results using a $\mathrm{KL}$-contamination model. If the distribution to be learnt is an unknown $P$ promised to belong to a class $\mathcal{C}$, the contaminated distribution $Q$ is some distribution satisfying $\mathsf{D_{KL}}(Q\|P)\leq\eta$. The noise is again non-additive, but the model is weaker than TV-contamination. Any $(\eta,A)$ approximation for $Q$ with respect to $\mathcal{C}$ yields a distribution $\widehat{P}$ such that $\mathsf{D_{KL}}(Q\|\widehat{P})\leq(A+1)\eta$. Therefore, we get that for Hellinger distance:
$\mathsf{H}(P,\widehat{P})\leq\mathsf{H}(P,Q)+\mathsf{H}(\widehat{P},Q)\leq\sqrt{\eta}+\sqrt{(A+1)\eta}\leq\sqrt{(2A+3)\eta}.$
Similarly, one can also bound $\mathsf{d_{TV}}(P,\widehat{P})=O(\sqrt{\eta})$ for constant $A$. To the best of our knowledge, the KL-contamination model has not been explicitly considered before, but if one were to directly apply the results of [15] with the assumption that $\mathsf{D_{KL}}(Q\|P)\leq\eta$, one would obtain a distribution $\widehat{P}$ such that $\mathsf{d_{TV}}(P,\widehat{P})=O(\sqrt{\eta\log 1/\eta})$, worse than ours by a $\sqrt{\log 1/\eta}$ factor which seems unavoidable using their approach [35]. Moreover, their results require that $\mathcal{C}$ be a class of balanced Bayes nets, a technical condition which is not needed for our analysis  [^2] A Bayes net is said to be $c$-balanced for some $c>0$ if all coordinates of the corresponding conditional probability table are between $c$ and $1-c$..
1.3.1 Open Problems
Our work opens up several interesting research avenues.
  • An intriguing question is whether we can extend our result for chordal graphs of bounded indegree to general graphs of bounded treewidth and bounded indegree. Interestingly, Stanley [71] showed that counting the number of acyclic orientations reduces to the evaluation of the Tutte polynomial at the point (2,0), and the Tutte polynomial can be evaluated efficiently for bounded treewidth graphs [61, 3]. This is relevant because the weights that EWA/RWM maintain are in some sense a weighted count of the number of acyclic orientations of the skeleton. However, we did not find a deletion-contraction recurrence for these weights, and so their connection to the Tutte polynomial is unclear.
  • Another important follow-up direction for learning Bayes nets would be to search over Markov equivalence classes rather than DAG’s. A Markov equivalence class corresponds to the set of DAGs that represent the same class of Bayes nets, and they can be represented as partially directed graphs (essential graphs) that satisfy some special graphical properties. It would be interesting to explore if the structure of essential graphs can be used to speed up weighted counting and sampling; indeed, a very recent work [68] gives a polynomial time algorithm for uniformly sampling an essential graph that is consistent with a given skeleton.
  • What is the role of approximate sampling in the context of distribution learning? So far, in this work, we have only used exact sampling algorithms for spanning arborescences and acyclic orientations. Can Markov chain techniques be brought to good use here? Our work further motivates settling the complexity status of approximately counting the number of acyclic orientations of an undirected graph; this question is a long-standing open problem in the counting/sampling literature.
  • Finally, while we have restricted ourselves to learning Bayes nets here, our framework is quite general and also applies to learning other classes of distributions, such as Ising models and factor models. We leave these questions for future work.
Organization of the paper
The rest of the paper is organized as follows. In Section 2, we present the preliminaries required for this work. Section 3 establishes the connection between regret in online learning to KL divergence in the scenario of agnostic learning of distributions. It also presents several necessary techniques from online learning along with the EWA and RWM algorithms that will be used later in our work. In Section 4, we discuss our results on learning tree-structured distributions, and present our alternative proper learning algorithm in Appendix C. In Section 5, we present our results on learning chordal-structured distributions and in Appendix A, we give the lower bound of learning tree-structured distributions. Finally, in Appendix B, we design efficient learning algorithms for Bayes nets over graphs with bounded vertex cover.
2 Preliminaries
For integers $0<m\leq n$, let $[n]$ denote the set $\{1,\dots,n\}$, and let $[m,n]$ denote the set $\{m,m+1,\dots,n\}$. For any $\eta>0$, let $\exp_{\eta}(u)$ denote $e^{-\eta u}$. For concise expressions and readability, we use the asymptotic complexity notion of $\widetilde{O}(\cdot)$, where we hide poly-logarithmic dependencies of the parameters. By stating i.i.d samples from a distribution $P$, we mean independently and identically distributed samples from $P$. For a positive integer $\ell$, $\mathsf{Unif}([\ell])$ denotes the uniform distribution on the set $[\ell]$, where each $i\in[\ell]$ is chosen with equal probability of $1/\ell$.
2.1 Probability Distributions
We let $\Delta(\mathcal{D})$ denote the set of probability distributions over the elements of a set $\mathcal{D}$. Let $P$ and $Q$ be two such distributions over $\mathcal{D}$:
  • The $\mathrm{KL}$-divergence between $P$ and $Q$ is defined as: $\mathsf{D_{KL}}(P,Q)=\sum_{x\in\mathcal{D}}P(x)\log\frac{P(x)}{Q(x)}$.
  • The total variation (TV) distance between $P$ and $Q$ is defined as: $\mathsf{d_{TV}}(P,Q)=\sum_{x\in\mathcal{D}}|P(x)-Q(x)|.$
Lemma 2.1 (Pinkser’s inequality).
Let $P$ and $Q$ be two probability distributions defined over the same sample space $\mathcal{D}$. Then the following holds:
$\mathsf{d_{TV}}(P,Q)\leq\sqrt{\frac{\mathsf{D_{KL}}(P||Q)}{2}}$
We study distributions over high-dimensional spaces, which can require an exponential amount of space to represent in general. So, for computational efficiency, we focus on distributions from which we can generate samples efficiently. This notion is formally defined below.
Definition 2.2 (Efficiently samplable distribution).
A distribution $P$ is said to be efficiently samplable if there exists a probabilistic Turing machine $M$ which on input $0^{k}$ produces a string $x$ such that $|\Pr[M(0^{k})=x]-P(x)|\leq 2^{-k}$ and $M$ runs in time $\mathrm{poly}(|x|+k)$.
Next we define graphical models of interest in this work. For a directed graph $G$ on $n$ nodes, we will identify its set of nodes with $[n]$. The underlying undirected graph of $G$ is called the skeleton. For any $i\in[n]$, let $\mathsf{pa}_{G}(i)$ denote the set of the parents of node $i$ and $\mathsf{nd}_{G}(i)$ denote the set of its non-descendants. The subscript $G$ may be removed if it is clear from the context.
Let us start with the definition of Bayesian networks.
Definition 2.3 (Bayesian networks).
A probability distribution $P$ over $n$ variables $X_{1},\dots,X_{n}$ is said to be a Bayesian network (Bayes net in short) on a directed acyclic graph $G$ with $n$ nodes if [^3] We use the notation $X_{S}$ to denote $\{X_{i}:i\in S\}$ for a set $S\subseteq[n]$. for every $i\in[n]$, $X_{i}$ is conditionally independent of $X_{\mathsf{nd}(i)}$ given $X_{\mathsf{pa}(i)}$. Equivalently, $P$ admits the factorization:
$P(x)=\Pr_{X\sim P}[X=x]=\prod_{i=1}^{n}\Pr_{X\sim P}[X_{i}=x_{i}\mid\forall j\in\mathsf{pa}(i),X_{j}=x_{j}]\qquad\text{for all }x.$ (1)
It is well-known that Bayesian networks are efficiently sampleable.
Now we define tree-structured distributions, a subclass of Bayesian networks defined above.
Definition 2.4 (Tree-structured distribution).
Let $T$ be a tree, and $G$ be any rooted orientation of $T$. A probability distribution $P$ is said to be $T$-structured if it is a Bayesian network on $G$. A distribution $P$ is said to be tree-structured if it is $T$-structured for some tree $T$.
Now we define the notion of polytree-structured distributions, which generalizes tree-structured distributions defined before.
Definition 2.5 (Polytree-structured distribution).
A directed acyclic graph (DAG) $G$ is said to be a polytree if the skeleton of $G$ is a forest. For a positive integer $k\in\mathbb{N}$, $G$ is said to be a $k$-polytree if the in-degree of each node in $G$ is at most $k$. A distribution $P$ is said to be a ($k$-)polytree-structured if it is a Bayes net defined over some ($k$-)polytree $G$.
Finally, we define the notion of Chordal-structured distributions, a class of Bayes nets that generalizes polytree-structured distributions and will be crucially used in this work.
Definition 2.6 (Chordal-structured distribution).
An undirected graph $G$ is said to be chordal if every cycle of length at least $4$ has a chord, that is, an edge that connects two non-adjacent vertices on the cycle. A distribution $P$ is said to be chordal-structured if it is a Bayes net defined over a DAG whose skeleton is chordal.
Clearly, any tree-structured distribution is polytree-structured, and any polytree-structured distribution is chordal-structured.
2.2 Clipping and Bucketing
We define two operations on probability distributions that will be useful in what follows. Let $P\in\Delta([k])$ for some positive integer $k$, and without loss of generality $P(1)\leq P(2)\leq\cdots\leq P(k)$. Note that $P(k-1)\leq 1/2$.
Now we define the notion of Clipping below.
Definition 2.7.
Given $0<\tau<1/k^{2}$, if $P(i)\geq\tau$ for $1\leq i<k$, define $\mathsf{Clip}_{\tau}(P)=P$. Otherwise, let $m$ such that $P(m)<\tau$ but $P(m+1)\geq\tau$, and define $\mathsf{Clip}_{\tau}(P)=Q$ where:
$Q(i)=\begin{cases}\tau&\text{if }1\leq i\leq m\\ P(i)&\text{if }m+1\leq i<k\\ 1-\sum_{j=1}^{k-1}Q(j)&\text{if }i=k\end{cases}$
Note that $\mathsf{Clip}_{\tau}(P)(k)=P(k)-\sum_{i=1}^{m}(\tau-P(i))\geq\frac{1}{k}-(k-1)\tau\geq\tau$ for $\tau\leq\frac{1}{2k^{2}}$.
Lemma 2.8.
For any $P^{*},P\in\Delta([k])$, if $\tau\leq 1/2k^{2}$ and $Q=\mathsf{Clip}_{\tau}(P)$:
$\mathsf{D_{KL}}(P^{*}\|Q)\leq\mathsf{D_{KL}}(P^{*}\|P)+2k^{2}\tau.$
Proof.
By definition, $Q(i)\geq P(i)$ for $i<k$ and $Q(k)\geq P(k)-k\tau$. Hence,
$\mathsf{D_{KL}}(P^{*}\|\mathsf{Clip}_{\tau}(P))-\mathsf{D_{KL}}(P^{*}\|P)$ $=\sum_{i=1}^{k}P^{*}(i)\log\frac{P(i)}{Q(i)}\leq\log\frac{P(k)}{Q(k)}\leq\log\frac{P(k)}{P(k)-k\tau}\leq\log\frac{1/k}{1/k-k\tau}$
$\leq 2k^{2}\tau$
where we used the facts that $P(k)\geq 1/k$, $-\log(1-x)\leq 2x$ for $x\leq 1/2$, and $k^{2}\tau\leq 1/2$. ∎
Now we define the bucketing operation. Below, let $Q=\mathsf{Clip}_{\tau}(P)$ so that $\tau\leq Q(1)\leq Q(2)\leq\cdots\leq Q(k-1)\leq 1/2$.
Definition 2.9.
Given $Q$ as above, and a parameter $\gamma>0$, define $\mathsf{Bucket}_{\gamma}(Q)$ to be a distribution $R$ such that:
$R(i)=\begin{cases}\tau\cdot(1+\gamma)^{\lfloor\log_{1+\gamma}(Q(i)/\tau)\rfloor}&\text{if }1\leq i<k\\ 1-\sum_{j=1}^{k-1}R(j)&\text{if }i=k\end{cases}$
Lemma 2.10.
Suppose $0<\tau<1/2k^{2}$ and $10\leq\gamma<$ 1. For any $P^{*},P\in\Delta([k])$, if $Q=\mathsf{Clip}_{\tau}(P)$, and $R=\mathsf{Bucket}_{\gamma}(Q)$, then:
$\mathsf{D_{KL}}(P^{*}\|R)\leq\mathsf{D_{KL}}(P^{*}\|Q)+\gamma\leq\mathsf{D_{KL}}(P^{*}\|P)+2k^{2}\tau+\gamma.$
Proof.
The second inequality follows from 2.8; so we focus on the first. Note that $R(i)\leq Q(i)$ for $i<k$, and so, $R(k)\geq Q(k)$.
$\mathsf{D_{KL}}(P^{*}\|R)-\mathsf{D_{KL}}(P^{*}\|Q)=\sum_{i=1}^{k}P^{*}(i)\log\frac{Q(i)}{R(i)}\leq\sum_{i=1}^{k-1}P^{*}(i)\log\frac{Q(i)}{R(i)}\leq\log(1+\gamma)\leq\gamma.$
For future reference, note that using the notation above, $R(k)\geq Q(k)\geq\tau$ (for $\tau\leq\frac{1}{2k^{2}}$).
2.3 PAC Distribution Learning
A distribution learning algorithm takes as input a sequence of i.i.d. samples generated from a probability distribution $P$ and outputs a description $\Theta$ of a distribution $\widehat{P}_{\Theta}$ as an estimate for $P$. In the following, $\mathcal{C}$ is a family of probability distributions over $[k]^{n}$, and $P$ is a distribution over $[k]^{n}$ not necessarily in $\mathcal{C}$. Let us first define the notion of approximation of a distribution that will be used in this work.
Definition 2.11 ($(\varepsilon,A)$-approximation of a distribution).
A distribution $\widehat{P}$ is said to be an $(\varepsilon,A)$-approximation for $P$ with respect to $\mathcal{C}$ if:
$\mathsf{D_{KL}}(P\|\widehat{P})\leq A\cdot\inf_{Q\in\mathcal{C}}\mathsf{D_{KL}}(P\|Q)+\varepsilon.$
When $A=1$, $\widehat{P}$ is said to be an $\varepsilon$-approximation for $P$. Note that when $P\in\mathcal{C}$, if $\widehat{P}$ is an $\varepsilon$-approximation for $P$, then $\mathsf{d_{TV}}(P,\widehat{P})\leq\sqrt{2\varepsilon}$ using Pinsker’s inequality (2.1).
Now we proceed to define the notion of PAC-learning with respect to KL divergence.
Definition 2.12 (PAC-learning in KL divergence).
A distribution learning algorithm is said to [^4] In the learning theory literature, when $A>1$, such a guarantee is also sometimes called semi-agnostic learning. be an agnostic PAC-learner for $\mathcal{C}$ with sample complexity $m_{\mathcal{C}}(n,k,\varepsilon,\delta)$ and running time $t_{\mathcal{C}}(n,k,\varepsilon,\delta)$, if for all distributions $P$ over $[k]^{n}$ and all $\varepsilon,\delta\in(0,1)$, given $\varepsilon,\delta$, and a sample set of size $m=m_{\mathcal{C}}(n,k,\varepsilon,\delta)$ drawn i.i.d. from $P$, the algorithm runs for time $t\leq t_{\mathcal{C}}(n,k,\varepsilon,\delta)$ and outputs the description $\Theta$ of a distribution $\widehat{P}_{\Theta}$ such that with probability at least $1-\delta$, $\widehat{P}_{\Theta}$ is an $\varepsilon$-approximation for $P$ with respect to $\mathcal{C}$ (where the probability is taken over the samples as well as the algorithm’s randomness).
If $P$ is restricted to be in $\mathcal{C}$, the algorithm is said to be a realizable PAC-learner for $\mathcal{C}$. If the output $\widehat{P}_{\Theta}$ is guaranteed to be in $\mathcal{C}$, the algorithm is said to be a proper PAC-learner for $\mathcal{C}$; otherwise the learner is called an improper PAC-learner.
2.4 Online Learning
The framework of prediction with experts is setup as follows; the formulation we follow here is based on Chapter 2 of [13]. The goal is to design an algorithm $\mathcal{A}$ that predicts an unknown sequence $x^{(1)},x^{(2)},\dots$ of elements of an outcome space $\mathcal{X}$. The algorithm’s predictions $\widehat{P}_{1},\widehat{P}_{2},\dots$ belong to a decision space $\mathcal{D}$ (which in our case will be the probability simplex on $\mathcal{X}$). The algorithm $\mathcal{A}$ makes its predictions sequentially, and the quality of its predictions is benchmarked against a set of reference predictions called experts. At each time instant $t$, $\mathcal{A}$ has access to the set of expert predictions $\mathcal{E}$, where $\mathcal{E}$ is a fixed subset of $\mathcal{D}$. $\mathcal{A}$ makes its own prediction $\widehat{P}_{t}$ based on the expert predictions. Finally, after the prediction $\widehat{P}_{t}$ is made, the true outcome $x^{(t)}$ is revealed, and the algorithm $\mathcal{A}$ incurs a loss $\ell(\widehat{P}_{t},x^{(t)})$, where $\ell:\mathcal{D}\times\mathcal{X}\to\mathbb{R}$ is a non-negative loss function.
Note that $x^{(1)},x^{(2)},\dots$ form an arbitrary sequence. Hence, naturally, the loss incurred by the algorithm depends on how well the experts fit these outcomes.
Now we are ready to define the notion of regret.
Definition 2.13 (Regret).
Given the notation above, the cumulative regret (or simply regret) over $T$ steps is defined as:
$\mathsf{Reg}_{T}(\mathcal{A};\mathcal{E})=\sum_{t=1}^{T}\ell(\widehat{P}_{t},x^{(t)})-\min_{E\in\mathcal{E}}\sum_{t=1}^{T}\ell(E,x^{(t)})$
Since $\mathcal{E}$ is usually fixed, we often use the shorthand $\mathsf{Reg}_{T}(\mathcal{A})$. The average regret is defined as $\frac{1}{T}\mathsf{Reg}_{T}(\mathcal{A})$. The total $T$ steps are also sometimes referred to as the horizon of the algorithm.
Our work utilizes two standard online learning algorithms: Exponentially Weighted Average (EWA) and Randomized Weighted Majority (RWM). Both algorithms assume that the expert set is finite; the first algorithm is deterministic, while the second one is probabilistic.
Fact 2.14.
In both EWA and RWM algorithms, when run with a hyperparameter $\eta>0$ and a set of experts =$\mathcal{E}=\{E_{1},\ldots,E_{N}\}$, the weight of the $i$-th expert ($i\in[N]$) after time step $t$ ($t\in[T]$) is denoted by $w_{i,t}$ defined as follows:
$w_{i,t}\triangleq\exp\left(-\eta\sum_{s=1}^{t}\ell(E_{i},x^{(s)})\right).$
Input :   Experts =${\mathcal{E}}=\{E_{1},\ldots,E_{N}\}$, parameter $\eta$, horizon $T$.
1 $w_{i,0}\leftarrow 1\mbox{ for each }i\in[N].$
2 for $t\leftarrow 1$ to $T$ do
3       $\widehat{P}_{t}\leftarrow\frac{\sum_{i=1}^{N}w_{i,t-1}\,{E_{i}}}{\sum_{j=1}^{N}w_{j,t-1}}$.
4       Output $\widehat{P}_{t}$.
5       Observe outcome $x^{(t)}$.
6       for $i\in[N]$ do
7             $w_{i,t}\leftarrow w_{i,t-1}\cdot\exp\left(-\eta\cdot\ell({E_{i}},x^{(t)})\right)$.
Algorithm 1 EWA forecaster
Input :   Experts =${\mathcal{E}}=\{E_{1},\ldots,E_{N}\}$, parameter $\eta$, horizon $T$.
1 $w_{i,0}\leftarrow 1\mbox{ for each }i\in[N].$
2 for $t\leftarrow 1$ to $T$ do
3       Sample $\widehat{P}_{t}\in\mathcal{E}$ where $\Pr[\widehat{P}_{t}={E_{i}}]=\frac{w_{i,t-1}}{\sum_{j=1}^{N}w_{j,t-1}}$.
4       Output $\widehat{P}_{t}$.
5       Observe outcome $x^{(t)}$.
6       for $i\in[N]$ do
7             $w_{i,t}\leftarrow w_{i,t-1}\cdot\exp\left(-\eta\cdot\ell({E_{i}},x^{(t)})\right)$.
Algorithm 2 RWM forecaster
Before proceeding to give the regret bounds for EWA and RWM algorithms, we need the notion of exp-concave loss functions.
Definition 2.15 (Exp-concave loss function).
Let $\mathcal{D}$ be a convex decision space, $\mathcal{X}$ be a sample space, and let $\alpha\in\mathbb{R}$ be a parameter. A loss function $\ell:\mathcal{D}\times\mathcal{X}\rightarrow\mathbb{R}$ is said to be $\alpha$-exp-concave if the function $f_{x}:\mathcal{D}\rightarrow\mathbb{R}$, $f_{x}(P)=\exp(-\alpha\ell(P,x))$ is concave for all $x\in\mathcal{X}$.
The following regret bound holds for EWA with exp-concave loss functions. Importantly for us, $\ell(P,x)=-\log(P(x))$ is exp-concave for $0\leq\alpha\leq 1$.
Lemma 2.16 ([13], Theorem 3.2).
With $N=|{\mathcal{E}}|$ experts and an $\alpha$-exp-concave loss function $\ell(P,x)$, for any sequence of outcomes $x^{(1)},\dots,x^{(T)}$, the regret of the EWA forecaster with the parameter $\eta=\alpha$ is bounded as follows:
$\mathsf{Reg}_{T}(\mathsf{EWA};\mathcal{E})\leq\frac{\log N}{\eta}.$
The regret bound below for RWM assumes the loss function is bounded.
Lemma 2.17 ([13], Lemma 4.2).
With $N=|\mathcal{E}|$ experts and a loss function bounded in $[-L_{\max},L_{\max}]$, for any sequence of outcomes $x^{(1)},\dots,x^{(T)}$, the expected regret of the RWM forecaster with the parameter $\eta=\sqrt{8(\log N)/T}$ is bounded as follows:
$\mathop{\mathbb{E}}[\mathsf{Reg}_{T}(\mathsf{RWM};\mathcal{E})]\leq L_{\max}\sqrt{\frac{T\log N}{2}}.$
Moreover, with probability at least $1-\delta$, the following holds:
$\mathsf{Reg}_{T}(\mathsf{RWM};\mathcal{E})]\leq L_{\max}\left(\sqrt{\frac{T\log N}{2}}+\sqrt{\frac{T}{2}\log\frac{1}{\delta}}\right).$
3 Agnostic Learning of Bayesian Networks
In this section, we study the notions and our results on the agnostic learning of Bayesian networks. We start with describing the relation between regret and KL divergence, which is crucially used throughout this work.
3.1 Connection between Regret and KL divergence
In the prediction with experts framework, let us fix the decision space $\mathcal{D}$ to be the probability simplex on the outcome space $\mathcal{X}$ and the loss function to be $\ell(P,x)=-\log P(x)$ for the distribution $P$ defined over the probability simplex on $\mathcal{X}$. Then we have the following general connection between the expected average regret and the expected KL divergence when the outcomes $x^{(1)},x^{(2)},\cdots$ are i.i.d samples from a distribution $P^{\ast}$.
Lemma 3.1.
If $\mathcal{A}$ is a (possibly randomized) online learning algorithm that receives i.i.d samples $x^{(1)},\ldots,x^{(T)}$ from a probability distribution $P^{*}$ on $\mathcal{X}$ and makes predictions $\widehat{P}_{1},\dots,\widehat{P}_{T}\in\mathcal{D}\subseteq\Delta(\mathcal{X})$ with respect to an expert set $\mathcal{C}\subseteq\mathcal{D}$,
$\mathop{\mathbb{E}}\limits_{x^{(1)},\ldots,x^{(T)}}\mathop{\mathbb{E}}_{t\sim\mathsf{Unif}([T])}\left[\mathsf{D_{KL}}(P^{*}\|\widehat{P}_{t})\right]\leq\frac{1}{T}\mathop{\mathbb{E}}\limits_{x^{(1)},\ldots,x^{(T)}}\left[\mathsf{Reg}_{T}(\mathcal{A}\,;\,\mathcal{C})\right]+\min\limits_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P).$
Note that if $\mathcal{A}$ is a randomized algorithm, both the left and right hand side of the inequality in 3.1 are random variables.
Proof.
From the definition of regret (2.13), we have:
$\mathsf{Reg}_{T}(\mathcal{A}\,;\,\mathcal{C})\triangleq\sum_{t=1}^{T}\log\frac{1}{\widehat{P}_{t}(x^{(t)})}-\min_{P\in\mathcal{C}}\sum_{t=1}^{T}\log\frac{1}{P(x^{(t)})}.$
Thus, the expectation over the samples of the average regret for $T$ rounds would be the following:
$\frac{1}{T}\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T)}\sim P^{*}}\left[\mathsf{Reg}_{T}(\mathcal{A}\,;\,\mathcal{C})\right]$ $=$ $\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T)}}\frac{1}{T}\cdot\left(\sum_{t=1}^{T}\log\frac{1}{\widehat{P}_{t}(x^{(t)})}-\min_{P\in\mathcal{C}}\sum_{t=1}^{T}\log\frac{1}{P(x^{(t)})}\right)$ (2)
$\geq$ $\frac{1}{T}\mathop{\mathbb{E}}_{x^{(1)},\dots,x^{(T)}}\left[\sum_{t=1}^{T}\log\frac{1}{\widehat{P}_{t}(x^{(t)})}\right]-\min_{P\in\mathcal{C}}\mathop{\mathbb{E}}_{x\sim P^{*}}\left[\log\frac{1}{P(x)}\right]$
$=$ $\frac{1}{T}\mathop{\mathbb{E}}_{x^{(1)},\dots,x^{(T)}}\left[\sum_{t=1}^{T}\log\frac{1}{\widehat{P}_{t}(x^{(t)})}\right]-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)-H(P^{*})$
where the second-last step is by applying Jensen’s inequality, as $\min$ is a concave function, and also uses the fact that $x^{(1)},\ldots,x^{(T)}\sim P^{\ast}$. Also, $H(P^{*})$ denotes the entropy of the distribution $P^{*}$. Thus we have the following:
$\frac{1}{T}\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T)}}\left[\mathsf{Reg}_{T}(\mathcal{A}\,;\,\mathcal{C})\right]$$\geq$ $\frac{1}{T}\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T)}}\left[\sum_{t=1}^{T}\log\frac{1}{\widehat{P}_{t}(x^{(t)})}\right]-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)-H(P^{*})$
$=$ $\frac{1}{T}\sum_{t=1}^{T}\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(t-1)}}\mathop{\mathbb{E}}_{x\sim P^{*}}\left[\log\frac{1}{\widehat{P}_{t}(x)}\Big{|}x^{(1)},\ldots,x^{(t-1)}\right]-H(P^{*})$
$\qquad\qquad\qquad-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)$
$=$ $\frac{1}{T}\sum_{t}\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(t-1)}}\left[\mathsf{D_{KL}}(P^{*}\|\widehat{P}_{t})\right]-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)$
$=$ $\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T-1)}}\left[\frac{1}{T}\sum_{t=1}^{T}\mathsf{D_{KL}}(P^{*}\|\widehat{P}_{t})\right]-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)$
In the second line, inside the summation over $t$, the expectation does not include $x^{(t)}$ because the prediction $\widehat{P}_{t}$ does not depend upon $x^{(t)}$, as $\widehat{P}_{t}$ is predicted before $x^{(t)}$ is revealed . ∎
We immediately have the following corollary:
Corollary 3.2.
In the same setup as 3.1 above, if $\frac{1}{T}\mathsf{Reg}_{T}(\mathcal{A};\mathcal{C})\leq\rho$,
$\mathop{\mathbb{E}}\limits_{x^{(1)},\ldots,x^{(T)}}\left[\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)\right]\leq\rho+\min\limits_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P).$
Proof.
This follows from applying the fact that KL-divergence is convex in its second argument. ∎
Now, we apply these results to the EWA and RWM algorithms. Let us start with the results for the EWA algorithm.
Lemma 3.3.
In the same setup as 3.1 above, suppose that for every $P\in\mathcal{C}$, $\min_{x\in\mathcal{X}}P(x)\geq\tau$ holds.
  • If $\mathcal{A}$ is the EWA algorithm run with parameter $\eta=1$, then we have the following:
    $\mathop{\mathbb{E}}_{x^{(1)},\dots,x^{(T)}}\left[\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)\right]\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\frac{\log|\mathcal{C}|}{T}.$
  • For any $\varepsilon,\delta\in(0,1)$, let $\mathcal{A}$ be the EWA algorithm run with parameter $\eta=\frac{\varepsilon}{2\sqrt{T}\log(1/\tau)\sqrt{\log(1/\delta)}}$. Then with probability at least $1-\delta$, the following holds:
    $\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\frac{2\log|\mathcal{C}|\log(1/\tau)\sqrt{\log(1/\delta)}}{\varepsilon\sqrt{T}}+\varepsilon.$
Proof.
Recall that our loss function $\ell(P,x)=-\log P(x)$ is $\eta$-exp-concave for all $0\leq\eta\leq 1$. Hence, we can apply 2.16 for any such $\eta$. The first item follows from 3.2 by setting $\eta=1$.
For the second item, we use the method of bounded differences. Define the function $f$:
$f(X^{(1)},\dots,X^{(T)})=\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)$
Claim 3.4.
For any $X^{(1)},X^{(2)},\dots,X^{(T)},{X^{(1)}}^{\prime}\in\mathcal{X}$,
$|f(X^{(1)},X^{(2)},\dots,X^{(T)})-f({X^{(1)}}^{\prime},X^{(2)},\dots,X^{(T)})|\leq 2\eta\log\frac{1}{\tau}.$
Proof.
Let $\widehat{P}_{t}$ and $\widehat{P}_{t}$ be the outputs from the EWA algorithm after $t$ rounds using the samples $X^{(1)},X^{(2)},\dots,X^{(T)}$ and ${X^{(1)}}^{\prime},X^{(2)},\dots,X^{(T)}$ respectively. Then, $\widehat{P}_{t}=\sum_{Q\in\mathcal{C}}w_{t,Q}Q$ and $\widehat{P}^{\prime}_{t}=\sum_{Q\in\mathcal{C}}w_{t,Q}^{\prime}Q$.
By definition of the EWA algorithm and the choice of our loss function, $w_{t,Q}=\frac{\left(\prod_{s\leq t}Q(X^{(s)})\right)^{\eta}}{\sum_{R\in\mathcal{C}}\left(\prod_{s\leq t}R(X^{(s)})\right)^{\eta}}$ and $w^{\prime}_{t,Q}=\frac{\left(Q({X^{(1)}}^{\prime})\prod_{2\leq s\leq t}Q(X^{(s)})\right)^{\eta}}{\sum_{R\in\mathcal{C}}R({X^{(1)}}^{\prime})\left(\prod_{2\leq s\leq t}R(X^{(s)})\right)^{\eta}}$. Since for every $Q\in\mathcal{C}$, $\tau\leq\frac{Q(X^{(1)})}{Q({X^{(1)}}^{\prime})}\leq\frac{1}{\tau}$, $\tau^{2\eta}\leq\frac{w_{t,Q}}{w^{\prime}_{t,Q}}\leq\tau^{-2\eta}$ and hence for every $y\in\mathcal{X}$, $\tau^{2\eta}\leq\frac{\widehat{P}^{\prime}_{t}(y)}{\widehat{P}_{t}(y)}\leq\tau^{-2\eta}$. Therefore:
$\left|f({X^{(1)}}^{\prime},\dots,X^{(T)})-f(X^{(1)},\dots,X^{(T)})\right|=\left|\sum_{y\in\mathcal{X}}P^{*}(y)\log\left(\frac{\sum_{t=1}^{T}\widehat{P}^{\prime}_{t}(y)}{\sum_{t=1}^{T}\widehat{P}_{t}(y)}\right)\right|\leq 2\eta\log\frac{1}{\tau}.$
Applying the McDiarmid’s inequality,
$\Pr\left[\left|\mathsf{D_{KL}}(P^{*}||\widehat{P})-\mathop{\mathbb{E}}\left[\mathsf{D_{KL}}(P^{*}||\widehat{P})\right]\right|\geq\varepsilon\right]\leq 2\exp\left(-\frac{2\varepsilon^{2}}{4T\eta^{2}\log^{2}\tau^{-1}}\right)\leq\delta,$
by our choice of $\eta$. The second item then follows from 2.16, 3.2, and the above. ∎
Now let us discuss the guarantees of the RWM algorithm.
Lemma 3.5.
In the same setup as 3.1 above, suppose that $\mathcal{C}=\{P_{1},\ldots,P_{N}\}$ and that for every $P_{i}\in\mathcal{C}$, $\min_{x\in\mathcal{X}}P_{i}(x)\geq\tau$ holds.
  • If $\mathcal{A}$ is the RWM algorithm run with the parameter $\eta=\sqrt{(8\log N)/T}$ — giving proper predictions $\widehat{P}_{1},\ldots,\widehat{P}_{T}\in\mathcal{C}$ — and $\widehat{P}\leftarrow\widehat{P}_{t}\in\mathcal{C}$ where $t$ is sampled uniformly from $[T]$, then we have:
    $\mathop{\mathbb{E}}_{\mathcal{A},x^{(1)},\dots,x^{(T)}}\left[\mathsf{D_{KL}}\left(P^{*}\|\widehat{P}\right)\right]\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\frac{\log(1/\tau)\sqrt{T\log|\mathcal{C}|}}{T}.$
  • With the same algorithm $\mathcal{A}$ as above, for any $\delta\in(0,1)$, the following bound holds with probability $\geq 1-\delta$ over the randomness of RWM, the random choice of $t$ from $[T]$, and the randomness of the samples $x^{(1)},\ldots,x^{(T)}$:
    $\mathsf{D_{KL}}(P^{\ast}\|\widehat{P})\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{\ast}\|P)+\frac{\log(1/\tau)\left(\sqrt{\log|\mathcal{C}|}+\sqrt{\log(2/\delta)}\right)}{\delta\sqrt{T}}.$
Proof.
The first item follows from 3.1, taking the expectation of both sides over the randomness of $\mathcal{A}$, and the expected regret bound in 2.17. Note that we can take $L_{\max}=\log(1/\tau)$ since $P(x)\geq\tau$ for all $x\in\mathcal{X}$ and $P\in\mathcal{C}$, by the assumption of the lemma.
The RWM algorithm (see Algorithm 2) can equivalently be viewed as follows. Sample $U_{1},\ldots,U_{T}\sim\mathsf{Uniform}((0,1])$ independently. At each $t\in[T]$, predict $\widehat{P}_{t}\leftarrow P_{i_{t}}\in\mathcal{C}$, where $i_{t}\in[N]$ is the index such that $U_{t}\in\Bigg{(}\frac{\sum_{j=1}^{i_{t}-1}w_{j,t-1}}{\sum_{k=1}^{N}w_{k,t-1}},\frac{\sum_{j=1}^{i_{t}}w_{j,t-1}}{\sum_{k=1}^{N}w_{k,t-1}}\Bigg{]}$. Note that modulo the choice of $U=(U_{1},\ldots,U_{T})$, the RWM algorithm is deterministic.
From 2.17 (high probability bound), with probability $\geq 1-\delta/2$ over the random choice of $U$, $\mathrm{Regret}_{T}(\mathcal{A}\,;\,\mathcal{C})\leq\log(1/\tau)\left(\sqrt{\frac{T\log|\mathcal{C}|}{2}}+\sqrt{\frac{T}{2}\log\frac{2}{\delta}}\right)$. Conditioning on this event $\mathcal{E}$, applying 3.1, we get
$\mathop{\mathbb{E}}_{x^{(1)},\ldots,x^{(T)}}\mathop{\mathbb{E}}_{t\sim\mathsf{Unif}([T])}\left[\mathcal{Z}_{X,t}\triangleq\mathsf{D_{KL}}(P^{\ast}\|\widehat{P}_{t})-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{\ast}\|P)\,\middle|\mathcal{E}\right]\leq\frac{\log(1/\tau)\left(\sqrt{\log|\mathcal{C}|}+\sqrt{\log(2/\delta)}\right)}{2\sqrt{T}}$
Note that, since $\widehat{P}_{t}\in\mathcal{C}$ for all $t\in[T]$, $\mathcal{Z}_{X,t}\geq 0$ for all $x^{(1)},\ldots,x^{(T)}\in\mathcal{X}$ and $t\in[T]$. Thus, we can use Markov’s inequality to show that, conditioned on the event $\mathcal{E}$, with probability $\geq 1-\delta/2$ over the randomness of $x^{(1)},\ldots,x^{(T)}$ and the random choice of $t\sim\mathsf{Unif}([T])$,
$\mathsf{D_{KL}}(P^{\ast}\|\widehat{P})-\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{\ast}\|P)\leq\frac{\log(1/\tau)\left(\sqrt{\log|\mathcal{C}|}+\sqrt{\log(2/\delta)}\right)}{\delta\sqrt{T}}.$ (3)
Taking a union bound with event $\overline{\mathcal{E}}$ and the failure event of the above bound (3), we can see that (3) holds with probability $\geq 1-\delta$ over the randomness of algorithm $\mathcal{A}$, the random choice of $t$, and the randomness of the samples $x^{(1)},\ldots,x^{(T)}$. This completes the proof of the lemma. ∎
3.2 Discretization
We will typically want to design agnostic PAC-learners for classes of distributions $\mathcal{C}$ that are infinite, e.g., tree-structured distributions, etc. However, to apply the results of the previous section, we first need to finitize $\mathcal{C}$. In the current section, we show how to accomplish this when $\mathcal{C}$ is the class of Bayes nets over a fixed DAG $G$.
Definition 3.6.
Given $\varepsilon\in(0,1)$ and variables $Y$ and $X_{1},\dots,X_{d}$ over $[k]$, let us define $\mathcal{N}^{Y\mid X_{1},\dots,X_{d}}_{\varepsilon}$: the $\varepsilon$-discretization of $Y\mid\left(X_{1},\dots,X_{d}\right)$. Let $\tau=\frac{\varepsilon}{4k^{2}}$ and $\gamma=\frac{\varepsilon}{4}$. For each $\sigma\in[k]^{d}$, let $\mathcal{N}^{Y\mid(X_{1},\dots,X_{d})=\sigma}_{\varepsilon}$ equal the following family of distributions over $Y$:
$\bigcup_{j\in[k]}\left\{P_{\sigma}\in\Delta(Y)\mid P_{\sigma}(j)\geq\tau,\forall\ell_{1},\dots,\ell_{k}\in\left[0,\left\lfloor\log_{1+\gamma}\frac{1}{2\tau}\right\rfloor\right],\forall i\neq j,P_{\sigma}(i)=\tau(1+\gamma)^{\ell_{i}}\right\},$
and let $\mathcal{N}^{Y\mid X_{1},\dots,X_{d}}_{\varepsilon}$ be the Cartestian product of $\mathcal{N}^{Y\mid(X_{1},\dots,X_{d})=\sigma}_{\varepsilon}$ over all $\sigma\in[k]^{d}$.
Definition 3.7.
Let $G$ be a DAG with $n$ nodes. For an alphabet size $k\geq 2$ and parameter $\varepsilon\in(0,1)$, given variables $X_{1},\dots,X_{n}$, we define the following family of probability distributions over $(X_{1},\dots,X_{n})$:
$\mathcal{N}^{G}_{\varepsilon}=\prod_{i\in[n]}\mathcal{N}_{\varepsilon/n}^{X_{i}\mid X_{\mathsf{pa}(i)}}$
where the product is a Cartesian product. Each element of $\mathcal{N}^{G}_{\varepsilon}$ can identified as a Bayes net on $G$.
Now we have the following lemma which bounds the size of $\log\left|\mathcal{N}^{G}_{\varepsilon}\right|$.
Lemma 3.8.
Suppose $G$ has in-degree at most $d$. Then:
$\log\left|\mathcal{N}^{G}_{\varepsilon}\right|=O\left(nk^{d+1}\log\frac{nk}{\varepsilon}\right).$
Proof.
Using the fact that $\log(1+\gamma)\geq\gamma/2$ for $0\leq\gamma\leq 1$, we can say the following:
$\left|\mathcal{N}^{G}_{\varepsilon}\right|=\prod_{i=1}^{n}\prod_{\sigma\in[k]^{|\mathsf{pa}(i)|}}\left|\mathcal{N}_{\varepsilon/n}^{X_{i}\mid X_{\mathsf{pa}(i)}=\sigma}\right|\leq\left(\frac{8nk}{\varepsilon}\log\frac{8nk^{2}}{\varepsilon}\right)^{nk^{d+1}}\leq\left(\frac{64n^{2}k^{3}}{\varepsilon^{2}}\right)^{nk^{d+1}}.$
Taking $\log$ on the above expression, we have the result. ∎
Lemma 3.9.
Any $P\in\mathcal{N}^{G}_{\varepsilon}$ satisfies $\min_{x\in[k]^{n}}P(x)\geq\left(\frac{\varepsilon}{4nk^{2}}\right)^{n}$.
Proof.
Since the conditional probability distribution at each node satisfies that its minimum probability value is at least $\varepsilon/(4nk^{2})$ by definition of $\mathcal{N}^{G}_{\varepsilon}$, and the probability mass function of a Bayes net is defined as the product of the conditional probabilities at each node, the result follows. ∎
Lemma 3.10.
Let $\mathcal{C}^{G}\subseteq\Delta([k]^{n})$ be the set of all Bayes nets defined on a given DAG $G$ with $n$ nodes and maximum in-degree $d$. For any $P^{*}\in\Delta([k]^{n})$, the following holds:
$\min_{P\in\mathcal{N}^{G}_{\varepsilon}}\mathsf{D_{KL}}(P^{*}\|P)\leq\min_{P\in\mathcal{C}^{G}}\mathsf{D_{KL}}(P^{*}\|P)+\varepsilon.$
Proof.
Given the DAG $G$, for a distribution $Q\in\Delta([k]^{n})$, a node $i\in[n]$, and an assignment $\sigma\in[k]^{|\mathsf{pa}(i)|}$ to the parents of the node $i$, let $Q_{\mathsf{pa}(i)}\in\Delta([k]^{|\mathsf{pa}(i)|})$ be the marginal distribution of $Q$ on $\mathsf{pa}(i)$, and let $Q_{i,\sigma}\in\Delta([k])$ denote the conditional distribution of $Q$ on node $i$, conditioned on its parents having value $\sigma$. Note that here, $Q$ itself need not be a Bayes net.
Suppose $P\in\mathcal{C}^{G}$ minimizes $\mathsf{D_{KL}}(P^{*}\|P)$. Let $R_{i,\sigma}=\mathsf{Bucket}_{\gamma}\mathsf{Clip}_{\tau}(P_{i,a})$, where $\tau$ and $\gamma$ are as in 3.7, and let $R$ be the Bayes net defined by having the conditional distributions at each node be $R_{i,\sigma}$:
$\Pr_{X\sim R}[X=x]=\prod_{i=1}^{n}R_{i,x_{\mathsf{pa}(i)}}(x_{i})\qquad\text{for all }x.$
We first claim that $R$ belongs to $\mathcal{N}^{G}_{\varepsilon}$. This is because by the definition of the clipping and bucketing operations, each $R_{i,\sigma}$ satisfies that for all values $j$ in $[k]$ except one, $R_{i,\sigma}(j)$ is in the interval $[\tau,1/2]$ and equals $\tau$ times a non-negative power of $(1+\gamma)$ and at the exceptional $j$, $R_{i,\sigma}(j)\geq\tau$ (by the remark after 2.10). So, $\mathcal{N}^{G}_{\varepsilon}$ must contain $R$.
Using 2.10, $\mathsf{D_{KL}}(P^{*}_{i,\sigma}\|R_{i,\sigma})\leq\mathsf{D_{KL}}(P^{*}_{i,\sigma}\|P_{i,\sigma})+\varepsilon/n$. We now use the following fact:
Fact 3.11 (Lemma 3.3 in [8]).
Suppose $P,Q\in\Delta([k]^{n})$, and $Q$ is a Bayes net on a DAG $G$. Then the following holds:
$\mathsf{D_{KL}}(P\|Q)=J_{P,G}+\sum_{i\in[n]}\sum_{\sigma\in[k]^{|\mathsf{pa}(i)}}P_{\mathsf{pa}(i)}(\sigma)\cdot\mathsf{D_{KL}}(P_{i,\sigma}\|Q_{i,\sigma}),$ (4)
where $J_{P,G}$ is a constant that depends on $P$ and $G$ but not on $Q$.
Therefore, we can say that:
$\mathsf{D_{KL}}(P^{*}\|R)$ $=J_{P^{*},G}+\sum_{i\in[n]}\sum_{\sigma\in[k]^{|\mathsf{pa}(i)}}P^{*}_{\mathsf{pa}(i)}(\sigma)\cdot\mathsf{D_{KL}}(P^{*}_{i,\sigma}\|R_{i,\sigma})$
$\leq J_{P^{*},G}+\sum_{i\in[n]}\sum_{\sigma\in[k]^{|\mathsf{pa}(i)}}P^{*}_{\mathsf{pa}(i)}(\sigma)\cdot(\mathsf{D_{KL}}(P^{*}_{i,\sigma}\|Q_{i,\sigma})+\varepsilon/n)$
$=\mathsf{D_{KL}}(P^{*}\|Q)+\varepsilon.$
3.3 Sample Complexity for learning Bayes Nets
We can now collect the tools developed in Sections 3.1 and 3.2 to prove 1.1 from the introduction. The algorithm that yields the first part of 1.1 (improper learning) is quite simple to describe (Algorithm 3 below), corresponding to EWA forecaster with $\eta=1$. In the following theorem, we specify the $\mathcal{N}$ and $T$ to take so as to obtain desired guarantees.
Input :    ${\mathcal{N}}=\{P_{1},\ldots,P_{N}\},T$, hyperparameter $\eta>0$.
Output :    Sampler for $\widehat{P}$.
1 $w_{i,0}\leftarrow 1\mbox{ for each }i\in[N].$
2 for $t\leftarrow 1$ to $T$ do
3         $\mbox{Observe sample }x^{(t)}\sim P^{\ast}$.
4       for $i\in[N]$ do
5             $w_{i,t}\leftarrow w_{i,t-1}\cdot P_{i}(x^{(t)})^{\eta}$.
6 function EWA-Sampler()
7       Sample $\mbox{Sample }t\leftarrow[T]$ uniformly at random.
8       Sample $i\sim[N]$ with probability $\frac{w_{i,t-1}}{\sum_{j\in[N]}w_{j,t-1}}$.
9       return $x\sim P_{i}$.
10return EWA-Sampler
/* This is a sampler for $\widehat{P}$. */
Algorithm 3 EWA-based learning for Bayes nets
Input :    ${\mathcal{N}}=\{P_{1},\ldots,P_{N}\},T$, hyperparameter $\eta>0$.
Output :    $\widehat{P}\in\mathcal{N}$.
1 $w_{i,0}\leftarrow 1\mbox{ for each }i\in[N].$
2 for $t\leftarrow 1$ to $T$ do
3       Sample $i_{t}\mbox{ from }[N]$ with $\Pr(i_{t}=i)=\frac{w_{i,t-1}}{\sum_{j\in[N]}w_{j,t-1}}$.
4         $\mbox{Observe sample }x^{(t)}\sim P^{\ast}$.
5       for $i\in[N]$ do
6             $w_{i,t}\leftarrow w_{i,t-1}\cdot P_{i}(x^{(t)})^{\eta}$.
7Sample $t$ uniformly from $[T]$.
8 return $\widehat{P}\leftarrow P_{i_{t}}$.
[1em]
Algorithm 4 RWM-based learning for Bayes nets
Theorem 3.12 (Bayes net improper learning).
Let $\mathcal{C}$ be the class of distributions over $[k]^{n}$ that can be defined as Bayes nets over an unknown DAG with $n$ nodes and in-degree $d$.
  • There is an algorithm that for all $P^{*}\in\Delta([k]^{n})$ and all $\varepsilon\in(0,1)$, receives $m=O\left(\frac{nk^{d+1}}{\varepsilon}\log\frac{nk}{\varepsilon}\right)$ i.i.d. samples from $P^{*}$ and returns a distribution $\widehat{P}$ that is with probability at least $2/3$, an $(\varepsilon,3)$-approximation for $P^{*}$ with respect to $\mathcal{C}$.
  • There is a realizable PAC-learner for $\mathcal{C}$ with sample complexity $O\left(\frac{nk^{d+1}}{\varepsilon\delta}\log\frac{nk}{\varepsilon\delta}\right)$.
  • There is an agnostic PAC-learner for $\mathcal{C}$ with sample complexity $\widetilde{O}(n^{4}k^{2d+2}\varepsilon^{-4}\log(1/\delta))$.
Proof.
Let $\mathcal{G}$ be the collection of all DAGs on $n$ nodes that have in-degree $d$, then we can say that $|\mathcal{G}|\leq(n!)^{d+1}$. Let $\mathcal{N}=\bigcup\limits_{G\in\mathcal{G}}\mathcal{N}^{G}_{\varepsilon/6}$, using the notation of 3.7. Then, using 3.8, we have:
$\log|\mathcal{N}|=O\left(\log|\mathcal{G}|+nk^{d+1}\log\frac{nk}{\varepsilon}\right)=O\left(nk^{d+1}\log\frac{nk}{\varepsilon}\right).$ (5)
By 3.10 and the definition of $\mathcal{N}$,
$\min_{P\in\mathcal{N}}\mathsf{D_{KL}}(P^{*}\|P)\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\varepsilon/6.$ (6)
We begin with the first item. We run EWA-based learning (Algorithm 3) using $\mathcal{N}$ as the expert set, $\eta=1$, and loss function $\ell(P,x)=-\log P(x)$. By the first part of 3.3 and Markov’s inequality, with probability at least $2/3$:
$\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)\leq 3\left(\min_{P\in\mathcal{N}}\mathsf{D_{KL}}(P^{*}\|P)+\frac{\log|\mathcal{N}|}{T}\right).$ (7)
Using (5), (6) and our choice of $T$ now ensures that right-hand-side in (7) is at most $3\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\varepsilon$.
For the second item, let $\mathcal{N}^{\prime}=\bigcup\limits_{G\in\mathcal{G}}\mathcal{N}_{\varepsilon\delta/6}^{G}$. By 3.10, if $P^{*}\in\mathcal{C}$, $\min_{P\in\mathcal{N}}\mathsf{D_{KL}}(P^{*}\|P)\leq\varepsilon\delta/6$. We can now run the same argument as above with $\varepsilon$ being $\varepsilon\delta$ to obtain the result.
We now prove the third item. Suppose $\varepsilon,\delta\in(0,1)$. By 3.9, for every $P\in\mathcal{N}$, $\min_{x\in[k]^{n}}P(x)\geq\left(\frac{\varepsilon}{24nk^{2}}\right)^{n}$. We run EWA using $\mathcal{N}$ as the expert set, $\eta=\frac{4n\log|\mathcal{N}|\log(24nk^{2}/\varepsilon)\sqrt{\log(1/\delta)}}{\varepsilon\sqrt{T}}$ and the same loss function. By the second part of 3.3, we get that with probability at least $1-\delta$:
$\mathsf{D_{KL}}\left(P^{*}\|\frac{1}{T}\sum_{t=1}^{T}\widehat{P}_{t}\right)\leq\min_{P\in\mathcal{N}}\mathsf{D_{KL}}(P^{*}\|P)+\frac{4n\log|\mathcal{N}|\log(24nk^{2}/\varepsilon)\sqrt{\log(1/\delta)}}{\varepsilon\sqrt{T}}+\frac{\varepsilon}{2}$ (8)
Using (5) and (6), if $T=\Theta\left(n^{4}k^{2d+2}\varepsilon^{-4}\log^{4}(nk/\varepsilon)\log(1/\delta)\right)$, then the right-hand-side of (8) is at most $\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{*}\|P)+\varepsilon$.
The above theorem is interesting in the context of improper learning of Bayes nets with constant in-degree $d$, because it gives a nearly-optimal sample complexity (optimal up to logarithmic factors) for realizable PAC learning with constant success probability (second part) and for getting a $(\varepsilon,3)$-approximation with constant success probability (first part).
While the next theorem does not have significant advantages for general in-degree $d$ Bayesian networks due to the less efficient nature of Algorithm 4 and its higher sample complexity compared to established algorithms such as Chow-Liu for proper learning, we prove it to lay the groundwork for subclasses of Bayesian networks (such as trees, polytrees, and chordal graphs). For these subclasses, the sampling method employed in Algorithm 4 can be optimized to yield efficient algorithms for proper learning.
Theorem 3.13.
Let $\mathcal{C}$ be the class of distributions over $[k]^{n}$ that can be defined as Bayes nets over an unknown DAG with $n$ nodes and in-degree $d$. There is a proper agnostic $(\varepsilon,\delta)$-PAC-learner for $\mathcal{C}$ with sample complexity $\widetilde{O}\left(\frac{n^{3}k^{d+1}}{\delta^{2}\varepsilon^{2}}\log^{3}\left(\frac{nk}{\varepsilon}\right)\right)$ for any $\delta\in(0,1)$ and $\varepsilon>0$.
Proof.
Define $\mathcal{N}=\bigcup\limits_{G\in\mathcal{G}}\mathcal{N}^{G}_{\varepsilon/2}$ — where $\mathcal{G}$ is the collection of all DAGs on $n$ nodes with in-degree $d$ — the same way as in the proof of 3.12. We run the RWM-based learning algorithm (Algorithm 4) with $\mathcal{N}$ as the expert set, $\eta=\sqrt{(8\log|\mathcal{N}|)/T}$, and the log loss function. By the second part of 3.5, and noting that:
  1. (i)
    The maximum loss at any round is at most $L_{\max}=\log(1/\tau)=n\log(8nk^{2}/\varepsilon)$ (using 3.9).
  2. (ii)
    the size of the cover $\mathcal{N}$ satisfies $\log|\mathcal{N}|\leq O\left(nk^{d+1}\log\frac{nk}{\varepsilon}\right)$ (using 3.8 and the fact that $\log|\mathcal{G}|\leq(d+1)n\log n$).
we have that, for any $\delta>0$, with probability $\geq 1-\delta$ over the random samples as well as the randomness of the algorithm, we have:
$\mathsf{D_{KL}}(P^{\ast}\|\widehat{P})\leq\min_{P\in\mathcal{N}}\mathsf{D_{KL}}(P^{\ast}\|P)+\frac{n\log\left(\frac{8nk^{2}}{\varepsilon}\right)\left(\sqrt{C\cdot nk^{d+1}\log(nk/\varepsilon)}+\sqrt{\log(2/\delta)}\right)}{\delta\sqrt{T}},$
where $C$ is some constant $\geq 1$.
Using 3.10, the definition of $\mathcal{C}$, and the fact that $\mathcal{N}=\bigcup_{G\in\mathcal{G}}\mathcal{N}^{G}_{\varepsilon/2}$, we have that
$\mathsf{D_{KL}}(P^{\ast}\|\widehat{P})\leq\min_{P\in\mathcal{C}}\mathsf{D_{KL}}(P^{\ast}\|P)+\frac{n\log\left(\frac{8nk^{2}}{\varepsilon}\right)\left(\sqrt{C\cdot nk^{d+1}\log(nk/\varepsilon)}+\sqrt{\log(2/\delta)}\right)}{\delta\sqrt{T}}+\varepsilon/2.$
Now if we choose any $T\geq\frac{8Cn^{2}\log^{2}\left(\frac{8nk^{2}}{\varepsilon}\right)\left(nk^{d+1}\log(nk/\varepsilon)+\log(2/\delta)\right)}{\delta^{2}\varepsilon^{2}}$, we get $\mathsf{D_{KL}}(P^{\ast}\|\widehat{P})\leq\varepsilon/2+\varepsilon/2=\varepsilon$ with probability $\geq 1-\delta$ (this bound for $T$ uses the fact that $(\sqrt{x}+\sqrt{y})^{2}\leq 2(x+y)$ by the AM-GM inequality). This gives us the $\widetilde{O}\left(\frac{n^{3}k^{d+1}}{\delta^{2}\varepsilon^{2}}\log^{3}\left(\frac{nk}{\varepsilon}\right)\right)$ sample complexity for proper agnostic $(\varepsilon,\delta)$-PAC learning. ∎
4 Learning Tree-structured distributions
Let us start with the definition of arborescence, which will be used throughout this section.
Definition 4.1 (Arborescence).
A directed graph $G=(V,E)$ is an out-arborescence rooted at $v\in V$ if its skeleton (underlying undirected graph) is a tree (acyclic and connected) and there is a unique directed path from $v$ to $w$ for every $w\in V\setminus\{v\}$. An in-arborescence rooted at $v$ is similar, but with unique directed paths from all $w\in V\setminus\{v\}$ to $v$.
Given a vertex set $V=[n]$, let $\mathcal{G}_{\rm T}$ denote the set of out-arborescences rooted at node $1$.Note that each $G\in\mathcal{G}_{\rm T}$ will have $m=n-1$ edges. Any tree-structured distribution $P$ on $[k]^{n}$ is $G$-structured for some $G\in\mathcal{G}_{\rm T}$. Let $\mathcal{C}^{\rm\textsc{Tree}}$ denote the set of all tree-structured distributions on $[k]^{n}$. Now we have the following lemma.
Lemma 4.2.
If we consider the finite set $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}=\bigcup_{G\in\mathcal{G}_{\rm T}}\mathcal{N}^{G}_{\varepsilon}$ of distributions on $[k]^{n}$ (see 3.7), then for any distribution $P^{\ast}\in\Delta([k]^{n})$,
$\min_{Q\in\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}}\mathsf{D_{KL}}(P^{\ast}\|Q)\leq\min_{Q\in\mathcal{C}^{\rm\textsc{Tree}}}\mathsf{D_{KL}}(P^{\ast}\|Q)+\varepsilon,$
The size is bounded by $\log|\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}|\leq O(nk^{2}\log(nk/\varepsilon))$.
Proof.
From 3.8 and 3.10, for any $G\in\mathcal{G}_{\rm T}$, we can discretize the set of $G$-structured distributions to a finite set $\mathcal{N}^{G}_{\varepsilon}$ of size $\log|\mathcal{N}^{G}_{\varepsilon}|\leq O(nk^{2}\log(nk/\varepsilon))$ such that, for any distribution $P^{\ast}$ on $[k]^{n}$, $\min_{Q\in\mathcal{N}^{G}_{\varepsilon}}\mathsf{D_{KL}}(P^{\ast}\|Q)\leq\min\limits_{\mbox{$G$-structured distributions }Q}\mathsf{D_{KL}}(P^{\ast}\|Q)+\varepsilon$.
Thus, if we take $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}\leftarrow\bigcup_{G\in\mathcal{G}_{\rm T}}\mathcal{N}^{G}_{\varepsilon}$, then the following holds:
$\log|\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}|\leq\max_{G}\log|\mathcal{N}^{G}_{\varepsilon}|+\log|\mathcal{G}_{\rm T}|\leq O(nk^{2}\log(nk/\varepsilon))+O(n\log n)\leq O(nk^{2}\log(nk/\varepsilon)),$
and the guarantee on the minimum KL divergence from $P^{\ast}$ comes from the definition of tree-structured distributions and the construction of $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$. ∎
4.1 Sampling from the EWA/RWM Distribution
In this section, we describe our approach for sampling tree-structured distribution using EWA and RWM algorithms.
Following 3.7, each tree-structured distribution $P$ in $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$ can be factored as $(G,V_{P})$ where $G$ is the underlying rooted arborescence and $V_{P}=(p_{e}:e\in E(G),p_{1})$ is the vector of functions where $p_{e}(x_{i},x_{j})$ for $e=(i,j)$ corresponds to $\Pr_{X\sim P}(X_{j}=x_{j}\mid X_{i}=x_{i})$ and $p_{1}(x_{1})$ corresponds to $\Pr_{X\sim P}(X_{1}=x_{1})$, where $X_{1}$ denotes the root node of $G$. Using the notation of 3.6, $p_{1}\in\mathcal{N}^{X_{1}}_{\varepsilon/n}$ and for $e=(i,j)$, $p_{e}\in\mathcal{N}^{X_{j}\mid X_{i}}_{\varepsilon/n}$. For simplicity, we will call these sets $\mathcal{N}^{1}$ and $\mathcal{N}^{e}$ respectively in this section. Let us start with our result for EWA algorithm.
Lemma 4.3.
Let $\widehat{P}_{t}$ be the output of the EWA algorithm run with the expert set $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$, parameter $\eta$, and horizon $T\geq t$ for time step $t$. For any sequence of observed samples $x^{(1)},\dots,x^{(T)}\in[k]^{n}$, one can generate a sample from $\widehat{P}_{t}$ in polynomial time for bounded $k$.
Proof.
In the EWA algorithm, $\widehat{P}_{t}$ is a mixture of the distributions in $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$. For a distribution $P\in\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$ that factors into $(G,(p_{e}:e\in E(G),p_{1}))$ as above, let the weight of $P$ in $\widehat{P}_{t}$ be $\omega(P)$. We have the following:
$\omega(P)$
$=\frac{\prod_{s<t}\exp_{\eta}\log\left(p_{1}(x^{(s)}_{1})\cdot\prod_{e\in E(G)}p_{e}(x^{(s)}_{e})\right)}{\sum_{H\in\mathcal{G}_{\rm T}}\sum_{q_{e}\in\mathcal{N}^{e}:e\in E(H),q_{1}\in\mathcal{N}^{1}}\prod_{s<t}\exp_{\eta}\log\left(q_{1}(x^{(s)}_{1})\cdot\prod_{e\in E(G)}q_{e}(x^{(s)}_{e})\right)}$
$=\frac{\prod_{s<t}\exp_{\eta}\log p_{1}(x^{(s)}_{1})\cdot\prod_{e\in E(G)}\prod_{s<t}\exp_{\eta}\log p_{e}(x^{(s)}_{e})}{\sum_{q_{1}\in\mathcal{N}^{1}}\prod_{s<t}\exp_{\eta}\log{q_{1}(x^{(s)}_{1})}\cdot\sum_{H\in\mathcal{G}_{\rm T}}\prod_{e\in E(H)}\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}$
$=\frac{\prod_{s<t}\exp_{\eta}\log p_{1}(x^{(s)}_{1})}{\sum_{q_{1}\in\mathcal{N}^{1}}\prod_{s<t}\exp_{\eta}\log q_{1}(x^{(s)}_{1})}\cdot\frac{\prod_{e\in E(G)}\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}{\sum_{H\in\mathcal{G}_{\rm T}}\prod_{e\in E(H)}\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}\cdot$
$\qquad\qquad\prod_{e\in E(G)}\frac{\prod_{s<t}\exp_{\eta}\log p_{e}(x^{(s)}_{e})}{\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}.$ (9)
We can interpret (9) as follows: the first term gives the probability of sampling $p_{1}\in\mathcal{N}^{1}$, the second term gives the probability of sampling $G\in\mathcal{G}^{\rm T}$, and the third term gives the probability of sampling $p_{e}\in\mathcal{N}^{e}$ for each $e\in E(G)$.
  • $p_{1}$ is sampled from a product distribution over $\mathcal{N}^{1}$. So, it can be sampled in time $O(|\mathcal{N}^{1}|)\leq(nk/\varepsilon)^{O(k)}$.
  • $G$ is sampled from a weighted mixture of spanning arborescences of the complete directed graph on $n$ nodes, where the weight of a spanning arborescence $H$ is proportional to $\prod\limits_{e\in E(H)}\mathbf{w}(e)$ and $\mathbf{w}(e)=\sum\limits_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})$ for any $e=(i,j)\in[n]^{2}$. Sampling weighted arborescences is known to be in polynomial time [10, 16, 75, 37], using Tutte’s theorem; for completeness, we describe the algorithm in the Appendix C.
  • Each $p_{e}$ is sampled from a product distribution over $\mathcal{N}^{e}$. So, each of them can be sampled in time $O(|\mathcal{N}^{e}|)\leq(nk/\varepsilon)^{O(k^{2})}$.
Once $P$ is sampled, a sample from $P$ can be generated in polynomial time as $P$ is tree-structured and $k$ is bounded. This completes the proof of the lemma. ∎
Now let us state our result for RWM algorithm.
Lemma 4.4.
If the RWM algorithm is run with the expert set $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$, parameter $\eta$, and horizon $T$, the prediction $\widehat{P}_{t}$ at time step $t\leq T$ can be computed in polynomial time, for any sequence of observed samples $x^{(1)},\dots,x^{(T)}\in[k]^{n}$ and bounded $k$.
Proof.
This follows directly from the argument used to prove 4.3, since in RWM, $\widehat{P}_{t}$ is sampled during the execution of the learning algorithm rather than during the sampling stage as in EWA. ∎
4.2 Learning Guarantees
In this section, we will state the learning guarantees for tree-structured distributions for both the improper and proper settings. Let us start by describing our result for the improper setting.
Theorem 4.5.
Let $P^{*}$ be an unknown discrete distribution defined over $[k]^{n}$, and $\varepsilon\in(0,1)$ be a parameter. Given sample access to $P^{*}$, there exist algorithms outputting efficiently-sampleable distributions for each of the following guarantees:
  1. (i)
    for the output distribution $Q$, with probability at least $2/3$, $\mathsf{D_{KL}}(P^{*}\|Q)$ is an $(\varepsilon,3)$-approximation of $P^{*}$ and the sample complexity is $O\left(\frac{nk^{2}\log(nk/\varepsilon)}{\varepsilon}\right)$;
  2. (ii)
    the algorithm is a realizable PAC learner for $\mathcal{C}^{\textsc{Tree}}$ with sample complexity $O\left(\frac{nk^{2}\log(nk/\delta\varepsilon)}{\delta\varepsilon}\right)$.
  3. (iii)
    the algorithm is an agnostic PAC learner for $\mathcal{C}^{\textsc{Tree}}$ with sample complexity $O(n^{4}k^{4}\varepsilon^{-4}\log(1/\delta))$.
Proof.
The algorithm is the EWA-based improper learning algorithm (Algorithm 3), and the correctness and sample complexity follow from 3.12. The fact that the output distribution is efficiently sampleable is shown in 4.3. ∎
In the realizable case (where $P^{\ast}$ is a tree-structured distribution), this sample complexity bound matches the following lower-bound in [14, Appendix A.2] up to log-factors for constant error probability.
Lemma 4.6 (Restatement of lower bound result in Appendix A.2 of [14]).
Let $P^{\ast}$ be an unknown tree-structured distribution defined over $\{0,1\}^{n}$, and $\varepsilon\in(0,1)$ be a parameter. Any algorithm that outputs a distribution $Q$ which is an $\varepsilon$-approximation of $P^{\ast}$ with constant probability requires $\Omega(\frac{n}{\varepsilon})$ samples from $P^{\ast}$ [^5] The lower bound stated in [14] holds for general Bayes net distributions of indegree $\leq d$. For tree-structured distributions, we have $d=1$. Moreover, their lower bound is with respect to TV-distance, which translates to KL-divergence due to Pinkser’s inequality..
Now we will state the guarantees for proper learning of the tree-structured distributions.
Theorem 4.7.
Let $P^{*}$ be an unknown discrete distribution defined over $[k]^{n}$, and $\varepsilon>0,\delta\in(0,1)$ be parameters. Given sample access to $P^{*}$, there exists a proper, agnostic PAC-learner for $\mathcal{C}^{\textsc{Tree}}$ that takes $\widetilde{O}\left(\frac{n^{3}k^{2}}{\varepsilon^{2}\delta^{2}}\log^{3}\left(\frac{nk}{\varepsilon}\right)\right)$ samples from $P^{\ast}$ and runs in time $O(\mathsf{poly}(n,1/\varepsilon,1/\delta))$ for constant $k$.
Proof.
The algorithm is the RWM-based proper learning algorithm (Algorithm 4) applied with $\mathcal{N}=\mathcal{N}^{\textsc{Trees}}_{\varepsilon/2}$, $\eta=\sqrt{(8\log|\mathcal{N}|)/T}$ and the log loss function. The algorithm, by construction, will output a tree-structured distribution in $\mathcal{N}$. The correctness (agnostic PAC guarantee) and sample complexity follow from exactly the same argument as in 3.13, taking $d=1$ since tree-structured distributions are Bayes nets with in-degree $1$. The fact that the RWM-computation can be done in $\mathsf{poly}(n,k,\frac{1}{\varepsilon},\frac{1}{\delta})$ time, in spite of having exponentially many distributions in $\mathcal{N}$, follows from the fact that the sampling from the weight distribution, for each time step $t$, can be done efficiently when $\mathcal{N}=\mathcal{N}^{\textsc{Tree}}_{\varepsilon/2}$ (see 4.4). ∎
5 Learning Chordal-structured distributions
In this section, we show that chordal-structured distributions can be learnt efficiently. Before proceeding to describe our results, we first define some notions that will be used in our results and proofs.
5.1 Preliminaries about Chordal Graphs
Given an undirected graph $G=(V,E)$ and subsets $S,T\subseteq V$, we let $G[S]$ denote the induced subgraph of $G$ on $S$, $E(S)$ denote the edge set of $G[S]$, and $E(S,T)$ denote the set of edges with one endpoint in $S$ and the other in $T$. For a vertex $v$ of $G$, let $\mathsf{Nbr}_{G}(v)$ denote the set of adjacent vertices of $v$ in $G$ (vertices $u$ such that $\{u,v\}\in E(G)$).
An undirected graph is chordal if every cycle of length at least 4 contains a chord, that is an edge connecting two vertices of the cycle which is not part of the cycle.
Definition 5.1 (Clique tree).
Let $G=(V,E)$ be a graph with vertex set $V$ and edge set $E$. The clique tree, denoted by $\mathcal{T}^{G}$, of $G$ is a tree that has the maximal cliques of $G$ as its vertices and for every two maximal cliques $C$ and $C^{\prime}$, each clique on the path from $C$ to $C^{\prime}$ in $\mathcal{T}^{G}$ contains $C\cap C^{\prime}$. $\mathcal{T}^{G}$ has the induced subtree property: for every vertex $v\in V$, the set of nodes in $\mathcal{T}^{G}$ that contains $v$ forms a connected subtree of $\mathcal{T}^{G}$. The treewidth of $G$ equals one less than the size of the largest maximal clique in $G$.
Let $G$ be a chordal graph, and let $\mathcal{T}^{G}$ be a clique tree of $G$. Fix an arbitrary maximal clique $C_{r}\in V(\mathcal{T}^{G})$ and we view $\mathcal{T}^{G}$ as a tree rooted at $C_{r}$. We denote this rooted clique tree as $\mathcal{T}^{G}_{C_{r}}$ and as $\mathcal{T}_{C_{r}}$ if $G$ is clear from context. For any maximal clique $C$, let $\mathsf{Pa}_{\mathcal{T}_{C_{r}}}(C)$ denote the parent of $C$ in the rooted tree $\mathcal{T}_{C_{r}}$, and let $\mathcal{T}_{C}$ denote the subtree of $\mathcal{T}_{C_{r}}$ rooted at $C$. Let $V[\mathcal{T}_{C}]$ denote the vertex set of $G[\mathcal{T}_{C}]$, that is, $\bigcup_{C^{\prime}\in V(\mathcal{T}_{C})}C^{\prime}$. For notational convenience, we will use $\mathcal{T}_{C}$ to denote both the subtree of $\mathcal{T}^{G}$ as well the vertex set $V[\mathcal{T}_{C}]$ when the usage will be clear from the context. Thus, $G[\mathcal{T}_{C}]$ denotes the subgraph of $G$ induced by the the vertices in $V[\mathcal{T}_{C}]$.
Definition 5.2 (Separator set).
Let $G$ be a chordal graph, and let $\mathcal{T}_{C_{r}}$ be a rooted clique tree of $G$ for a maximal clique $C_{r}$. For $C\in V(\mathcal{T}_{C_{r}})$, the separator of $C$ with respect to $\mathcal{T}_{C_{r}}$ is defined as follows:
$\mathsf{Sep}(C)=C\cap\mathsf{Pa}_{\mathcal{T}_{C_{r}}}(C)$
Next we define the notion of link set, which is crucially used in our proofs.
Definition 5.3 (Link set).
Let $G=(V,E)$ be a chordal graph, and $\mathcal{T}_{C_{r}}$ be a rooted clique tree of $G$ for a maximal clique $C_{r}$. For $C\in V(\mathcal{T}_{C_{r}})$, the link set of $C$ is defined as follows:
$\mathsf{Link}(C)=E(C,V[\mathcal{T}_{C}])$
An orientation of an edge set $F$ assigns a direction to each edge in $F$; an orientation is acyclic if it does not give rise to a directed cycle. The indegree of an orientation is the maximum number of incident edges which are oriented inward at any vertex. Two acyclic orientations on edge sets $F$ and $F^{\prime}$ are said to be consistent with each other if they agree on the edges in $F\cap F^{\prime}$.
Definition 5.4 (Indegree-Bounded Acyclic Orientations).
Let $G=(V,E)$ be an undirected graph. For any subset of edges $F\subseteq E$ and integer $d\geq 0$, $\mathsf{AO}_{d}(F)$ denotes the set of all acyclic orientations of $F$ with indegree at most $d$. Given an orientation $\mathcal{O}$ that assigns orientations to a subset of $E$, we let $\mathsf{AO}_{d}(F;\mathcal{O})$ denote the subset of $\mathsf{AO}_{d}(F)$ that is consistent with $\mathcal{O}$. If $F$ corresponds to the edge set of a subgraph $H$, then we also use the notations $\mathsf{AO}_{d}(H)$ and $\mathsf{AO}_{d}(H,\mathcal{O})$ respectively.
{subcaptionblock}Refer to caption
Figure 2: A chordal graph and a clique tree decomposition with reference clique $C=\textrm{DEG}$. In the left panel, the edges of $\mathsf{Link}(C)$ are in red and the vertices $V[\mathcal{T}_{C}]$ are in bold. In the right panel, the nodes of $\mathcal{T}_{C}$ are in green, and the separator vertices in each node of the clique tree are colored in blue.
{subcaptionblock}Refer to caption
Figure 3: The same chordal graph and a different clique tree decomposition with reference clique $C=\textrm{ACDE}$. The colors have the same meaning as in (a).
We will use below some standard observations about chordal graphs and acyclic orientations; for the sake of completeness, we give their proofs.
Lemma 5.5 (Lemma 1 of [11], restated).
Let $G=(V,E)$ be a chordal graph and consider a clique tree $\mathcal{T}_{C_{r}}$ of $G$ rooted at a node $C_{r}$. Let $C$ be a node in $\mathcal{T}_{C_{r}}$ and $C_{1},\ldots,C_{\ell}$ be its children in $\mathcal{T}_{C_{\ell}}$. Then the edge sets of the graphs $G[\mathcal{T}_{C_{1}}\setminus\mathsf{Sep}(C_{1})],\ldots,G[\mathcal{T}_{C_{\ell}}\setminus\mathsf{Sep}(C_{\ell})]$ are mutually disjoint.
Proof.
We will prove this by contradiction. Let us assume that there exists $i\neq j$ with $i,j\in[\ell]$ such that $G[\mathcal{T}_{C_{i}}\setminus\mathsf{Sep}(C_{i})]$ and $G[\mathcal{T}_{C_{j}}\setminus\mathsf{Sep}(C_{j})]$ share an edge $e=\{u,v\}$. This implies that both $G[\mathcal{T}_{C_{i}}]$ and $G[\mathcal{T}_{C_{j}}]$ contain the edge $e$. However, this would imply $C\supseteq\{u,v\}$ by the inducted subtree property (since any pair of nodes in $\mathcal{T}_{C_{i}}$ and $\mathcal{T}_{C_{j}}$ can only be connected in $\mathcal{T}^{G}$ through $C$), and thus that $e$ will be present in both $E(\mathsf{Sep}(C_{i}))$ and $E(\mathsf{Sep}(C_{j}))$. This contradicts our assumption on $e$. ∎
Lemma 5.6 (See Theorem 11 in Chapter 4 of [72]).
Let $G$ be a chordal graph which is acyclically orientable with indegree $\leq d$, and consider a rooted clique tree $\mathcal{T}_{C_{r}}$ of $G$. Consider a non-leaf node $C\in\mathcal{T}_{C_{r}}$ and $C_{1},\ldots,C_{\ell}$ be the set of children of $C$ in the rooted tree. Consider an acyclic orientation $\mathcal{O}_{C}$ of $\mathsf{Link}(C)$ with indegree $\leq d$, and let $\mathcal{O}_{\mathsf{Sep}(C_{i})}$ be the orientation $\mathcal{O}_{C}$ restricted to $E(\mathsf{Sep}(C_{i}),\mathcal{T}_{C_{i}})$ for every $i\in[\ell]$. Then there exists a bijection between $\mathsf{AO}_{d}(\mathcal{T}_{C},\mathcal{O}_{C})$ and $\mathsf{AO}_{d}(\mathcal{T}_{C_{1}},\mathcal{O}_{\mathsf{Sep}(C_{1})})\times\ldots\times\mathsf{AO}_{d}(\mathcal{T}_{C_{\ell}},\mathcal{O}_{\mathsf{Sep}(C_{\ell})})$.
Proof.
Consider any acyclic orientation $\mathcal{O}$ of $\mathcal{T}_{C}$ with indegree $\leq d$ consistent with a given acyclic orientation $\mathcal{O}_{C}$ of $\mathsf{Link}(C)$. Any subset of an acyclic orientation $\mathcal{O}$ will be acyclic as well, and indegree bounds will be preserved. Hence $\mathcal{O}$ restricted to each $\mathcal{T}_{C_{i}}$ will give acyclic orientations $\mathcal{O}_{i}$ of $\mathcal{T}_{C_{i}}$ with indegree $\leq d$. Given these consistent orientations $\mathcal{O}_{1},\ldots,\mathcal{O}_{\ell}$ of the subtrees, and knowing that $\mathcal{O}$ is consistent with $\mathcal{O}_{C}$, we can reconstruct $\mathcal{O}$. This suffices to argue the injectivity.
In order to prove surjectivity, let us consider acyclic orientations $\mathcal{O}_{i}\in\mathsf{AO}_{d}(\mathcal{T}_{C_{i}},\mathcal{O}_{\mathsf{Sep}(C_{i})})$ for every $i\in[\ell]$. Now let us consider the following orientation $\mathcal{O}$ by taking the union of $\mathcal{O}_{C}$ and $\bigcup_{i=1}^{\ell}\mathcal{O}_{i}$. Following 5.5, we know that $\mathcal{O}$ is consistent with $\mathcal{O}_{C}$. Now we would like to prove that $\mathcal{O}$ is an acyclic orientation and has indegree $\leq d$.
We will prove acyclicity by contradiction. Assume that there is at least one cycle in $\mathcal{O}$. Consider the shortest cycle among all those cycles in $\mathcal{O}$. The cycle must contain at least two edges in $E(C)$, because if it contains one or no edges of $E(C)$, then it is contained inside some $\mathcal{T}(C_{i})$ which is impossible since $\mathcal{O}_{i}$ is acyclic. Consider two vertices $u,v\in C$ which are in the cycle but not adjacent in the cycle. In $C$, there exists an edge between $u$ and $v$, with $\mathcal{O}_{C}$ orienting it either $(u,v)$ or $(v,u)$. Whichever the case, we can shorten the cycle by taking a shortcut on the edge, which contradicts the assumption that we are considering the shortest cycle. This implies that $\mathcal{O}$ is acyclic.
To argue that the indegree of $\mathcal{O}$ is $\leq d$, note that any $v$ which is common to $\mathcal{T}_{C_{i}}$ and $\mathcal{T}_{C_{j}}$ for $i\neq j$ will belong to $C$ as well. Hence the orientation of all incident edges of $v$ will be fixed by $\mathcal{O}_{C}$ itself, and combining $\mathcal{O}_{1},\ldots,\mathcal{O}_{\ell}$ will not increase the indegree beyond $d$ since they are all consistent with $\mathcal{O}_{C}$. Thus the mapping is surjective as well. Combining the above, we have the proof of the lemma. ∎
Lemma 5.7.
Let $C$ and $C_{i}$ be two cliques of $G$ so that $C$ is a node in $\mathcal{T}^{G}$ and $C_{i}$ is a child of $C$ in $\mathcal{T}^{G}$. Then there are no edges between $V[\mathcal{T}_{C_{i}}]\setminus C$ and $C\setminus C_{i}$.
Proof.
Assume for the sake of contradiction that $\{u,v\}$ is an edge between $u\in V[\mathcal{T}_{C_{i}}]\setminus C$ and $v\in C\setminus C_{i}$. In particular, say $u\in K$ where $K$ is a clique in $\mathcal{T}_{C_{i}}$. Then, by the definition of the clique tree decomposition, (i) the edge is contained neither in $C$ nor in any of the cliques in the subtree rooted at $C_{i}$ (since $u\not\in C$ and $v\in C\setminus C_{i}$), but (ii) the edge is part of some maximal-clique $C^{\prime}$. Hence, $C^{\prime}$ and $K$ must be separated by $C$ in the clique tree. But this is a contradiction, since $u\in C^{\prime}$ and $u\in K$ but $u\not\in C$. ∎
5.2 EWA and RWM with Chordal Experts
In this section, we consider distributions on chordal graphs which can be oriented acyclically with indegree $\leq d$. This assumption is sufficient to bound the size of each maximal clique.
Remark 5.8.
Let $G$ be an undirected chordal graph. If there exists an acyclic orientation of $G$ with indegree $\leq d$, then for any clique tree decomposition $\mathcal{T}_{C_{r}}$ of $G$ (rooted at a maximal clique $C_{r}$), all the nodes of $\mathcal{T}_{C_{r}}$ (maximal cliques of $G$) have cardinality $\leq d+1$.
Proof.
By assumption, we have an acyclic orientation with indegree $\leq d$ for each clique $C\in V(\mathcal{T}_{C_{r}})$. If we topologically order the nodes of $C$, the last node will have indegree $|C|-1\leq d$, which implies $|C|\leq d+1$. ∎
We will establish the following theorem.
Theorem 5.9.
Let $G$ be an undirected chordal graph, and suppose $k$ and $d$ are fixed constants. Let $\mathcal{L}^{G}_{d}$ be the family of distributions over $[k]^{n}$ that can be defined as Bayes nets over DAGs having skeleton $G$ and with indegree $\leq d$. Given sample access from $P^{*}\in\Delta([k]^{n})$, and a parameter $\varepsilon>0$, there exist the following:
  • (i)
    An agnostic PAC-learner for $\mathcal{L}^{G}_{d}$ using $\widetilde{O}(n^{4}k^{2d+2}\varepsilon^{-4}\log(1/\delta))$ samples that is improper and returns an efficiently-samplable mixture of distributions from $\mathcal{L}^{G}_{d}$.
  • (ii)
    An agnostic PAC-learner $\mathsf{LearnChordalDist}$ for $\mathcal{L}^{G}_{d}$ using $\widetilde{O}(n^{3}k^{d+1}\varepsilon^{-2}\log(n/\varepsilon\delta))$ samples and $\mathrm{poly}(n)$ running time that is proper and returns a distribution from $\mathcal{L}^{G}_{d}$.
The sample complexity guarantees follow from 3.12 and from 3.13. What remains to be justified is that the algorithms run efficiently and the distribution output by the improper learning algorithm can be sampled efficiently. Below, let $x^{(1)},\dots,x^{(T)}$ be the samples drawn from $P^{*}$.
Let $\mathcal{T}^{G}_{C_{r}}$ be a clique tree decomposition of $G$ rooted at a maximal clique $C_{r}$. We can assume without loss of generality that $G$ has at least one acyclic orientation with indegree $\leq d$; otherwise, $\mathcal{L}^{G}_{d}$ is empty and the problem is trivial. Hence, by Remark 5.8, every node in $\mathcal{T}^{G}_{C_{r}}$ has at most $d+1$ vertices of $G$.
Claim 5.10.
For any $C\in V(\mathcal{T}_{C_{r}})$, $|\mathsf{AO}_{d}(\mathsf{Link}(C))|\leq\binom{n+d}{d}^{d+1}$.
Proof.
Note that there are at most $d+1$ vertices in $C$ and each such vertex has at most $n$ incident edges, of which at most $d$ can be incoming. Combining the above, we have the claim. ∎
For an orientation $\mathcal{O}$ of $G[\mathcal{T}_{C}]$ and a node $v\in V[\mathcal{T}_{C}]$, let $\mbox{in}_{C}(v,\mathcal{O})$ denote the set of in-neighbors of $v$ in $G[\mathcal{T}_{C}]$ with respect to $\mathcal{O}$. We define the weight, with respect to the clique $C$, of a node $v\in V[\mathcal{T}_{C}]$, acyclic orientation $\mathcal{O}$ of $G[\mathcal{T}_{C}]$ and time-step $t\in[T]$ as follows:
$\texttt{wt}_{C}(v,\mathcal{O},t)\triangleq\sum_{q\in\mathcal{N}_{\varepsilon/n}^{v\mid\text{in}_{C}(v,\mathcal{O})}}\prod_{s=1}^{t}\exp_{\eta}\log q\left(x^{(s)}_{v}\mid x^{(s)}_{\text{in}_{C}(v,\mathcal{O})}\right).$ (10)
For all $C\in V(\mathcal{T}^{G})$ and for all $\mathcal{O}_{C}\in\mathsf{AO}_{d}(\mathsf{Link}(C))$, and for all $t\in[T]$, we will store an entry in a table $\mathsf{Table}[C,\mathcal{O}_{C},t]$.
$\mathsf{Table}[C,\mathcal{O}_{C},t]\triangleq\sum_{\mathcal{O}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C}],\mathcal{O}_{C})}\prod_{v\in V[\mathcal{T}_{C}]}\texttt{wt}_{C}(v,\mathcal{O},t).$ (11)
Note that, according to this definition, $\mathsf{Table}[C_{r},\emptyset,t]$ gives the total weight of all indegree $\leq d$ acyclic orientations of $G$ after observing $t$ samples. The sum is over an exponential-sized set; nevertheless, we will be able to use dynamic programming to compute it efficiently.
Let us start with the case that $C$ is a leaf node of $\mathcal{T}^{G}$. Note that, if $C$ is a leaf node of $\mathcal{T}^{G}$, $\mathsf{AO}_{d}(G[\mathcal{T}_{C}],\mathcal{O}_{C})=\{\mathcal{O}_{C}\}$, where $\mathcal{O}_{C}$ is an acyclic orientation of $C$. So, the corresponding table entry can be directly computed for any $\mathcal{O}_{C}$ in $\mathsf{poly}(n,t,1/\varepsilon)$ time for constant $d$ and $k$. This completes the base case of our dynamic programming.
Now let us assume that $C$ has $\ell$ children $C_{1},\ldots,C_{\ell}$ in $\mathcal{T}^{G}$. For $\mathcal{O}_{C}\in\mathsf{AO}_{d}(\mathsf{Link}(C))$, we show that $\mathsf{Table}[C,\mathcal{O}_{C},t]$ can be inductively computed in terms of the table entries $\mathsf{Table}[C_{i},\mathcal{O}_{C_{i}},t]$ for $1\leq i\leq\ell$, where each $\mathcal{O}_{C_{i}}\in\mathsf{AO}_{d}(\mathsf{Link}({C_{i}}))$, so that we obtain a bottom-up dynamic programming algorithm for counting weighted acyclic orientations with indegree $\leq d$ starting from the leaf nodes of $\mathcal{T}^{G}$. By 5.6, we first break each acyclic orientation $\mathcal{O}$ of $\mathcal{T}_{C}$ into $\mathcal{O}_{1},\ldots,\mathcal{O}_{\ell}$ (acyclic orientations of the child subtrees $\mathcal{T}_{C_{i}}$) which can be combined consistently with an orientation $\mathcal{O}_{C}\in\mathsf{AO}_{d}(\mathsf{Link}(C))$.
The table entry $\mathsf{Table}[C,\mathcal{O}_{C},t]$ can be expressed as follows:
$\mathsf{Table}[C,\mathcal{O}_{C},t]$$\triangleq$ $\sum_{\mathcal{O}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C}],\mathcal{O}_{C})}\prod_{v\in V[\mathcal{T}_{C}]}\texttt{wt}_{C}(v,\mathcal{O},t)$ (12)
$=$ $\sum_{\begin{subarray}{c}\mathcal{O}_{1}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C_{1}}],\mathcal{O}_{C}),\\[-1.39998pt] \vdots\\ \mathcal{O}_{\ell}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C_{\ell}}],\mathcal{O}_{C})\end{subarray}}\prod_{v\in V[\mathcal{T}_{C}]}\texttt{wt}_{C}(v,\cup_{j=1}^{\ell}\mathcal{O}_{j},t)$
$=$ $\prod_{v\in C}\texttt{wt}_{C}(v,\mathcal{O}_{C},t)\sum_{\begin{subarray}{c}\mathcal{O}_{1}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C_{1}}],\mathcal{O}_{C}),\\[-1.39998pt] \vdots\\ \mathcal{O}_{\ell}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C_{\ell}}],\mathcal{O}_{C})\end{subarray}}\left(\prod_{i=1}^{\ell}\prod_{v\in V[\mathcal{T}_{C_{i}}]\setminus C}\texttt{wt}_{C_{i}}(v,\mathcal{O}_{i},t)\right),$
where in the last line, we used the fact that $V[\mathcal{T}_{C_{i}}]\setminus C$ are disjoint for distinct $i$, and also that, there are no edges between $V[\mathcal{T}_{C_{i}}]\setminus C$ and $C\setminus C_{i}$ by 5.7.
For each $i$, let $\mathsf{Cons}_{i}(\mathcal{O}_{c})\subseteq\mathsf{AO}_{d}(\mathsf{Link}({C_{i}}))$ be the set of acyclic orientations of $\mathsf{Link}(C_{i})$ of indegree at most $d$ that are consistent with $\mathcal{O}_{C}$. From (12), we can write $\mathsf{Table}[C,\mathcal{O}_{C},t]$ as:
$\prod_{v\in C\setminus(C_{1}\cup\ldots\cup C_{\ell})}\texttt{wt}_{C}(v,\mathcal{O}_{C},t)~{}\times$
$~{}~{}\sum_{\forall i,\mathcal{O}_{C_{i}}\in\mathsf{Cons}_{i}(\mathcal{O}_{C})}\sum_{\forall i,\mathcal{O}_{i}\in\mathsf{AO}_{d}(G[\mathcal{T}_{C_{i}}],\mathcal{O}_{C_{i}})}\left(\prod_{i=1}^{\ell}\prod_{v\in V[\mathcal{T}_{C_{i}}]\setminus\mathsf{Sep}(C_{i})}\texttt{wt}_{C_{i}}(v,\mathcal{O}_{i},t)\right)\cdot\left(\prod_{v\in\cup_{i=1}^{\ell}\mathsf{Sep}(C_{i})}\texttt{wt}_{C}(v,\mathcal{O}_{C},t)\right).$ (13)
We can now state a recurrence for $\mathsf{Table}[C,\mathcal{O}_{C},t]$ using the notation above.
Lemma 5.11.
For any $C\in V(\mathcal{T}_{C_{r}}),\mathcal{O}_{C}\in\mathsf{AO}_{d}(\mathsf{Link}(C)),t\in[T]$, if $C_{1},\dots,C_{\ell}$ are the children of $C$ in $\mathcal{T}_{C_{r}}$, then:
$\mathsf{Table}[C,\mathcal{O}_{C},t]=\prod_{v\in C\setminus\cup_{i}C_{i}}\mathsf{wt}_{C}(v,\mathcal{O}_{C},t)\sum_{\forall i,\mathcal{O}_{C_{i}}\in\mathsf{Cons}_{i}(\mathcal{O}_{C})}\left(\prod_{i=1}^{\ell}\mathsf{Table}[C_{i},\mathcal{O}_{C_{i}},t]\right)\Xi(\mathcal{O}_{C},\mathcal{O}_{C_{1}},\dots,\mathcal{O}_{C_{\ell}}),$
where:
$\Xi(\mathcal{O}_{C},\mathcal{O}_{C_{1}},\dots,\mathcal{O}_{C_{\ell}})=\prod_{v\in\cup_{i}\mathsf{Sep}(C_{i})}\frac{\mathsf{wt}_{C}(v,\mathcal{O}_{C},t)}{\prod\limits_{\begin{subarray}{c}j\in[\ell]:C_{j}\ni v\end{subarray}}\mathsf{wt}_{C_{j}}(v,\mathcal{O}_{C_{j}},t)}.$
Proof.
We argue by induction. For the base case, when $C$ is a leaf, $\mathsf{Table}[C,\mathcal{O}_{C},t]=\prod_{v\in C}\mathsf{wt}_{C}(v,\mathcal{O}_{C},t)$ which agrees with (11). Otherwise, we inductively use (11) on each $\mathsf{Table}[C_{i},\mathcal{O}_{C_{i}},t]$ and check that (13), and therefore (11), holds for $\mathsf{Table}[C,\mathcal{O}_{C},t]$. The only fact we need here is that if $v\in\mathsf{Sep}(C_{i})$, all its incident edges belong to $\mathsf{Link}(C_{i})$ and hence their orientations are fixed by $\mathcal{O}_{C_{i}}$. ∎
The pseudocode for counting the number of acyclic orientations with maximum indegree $d$ is described in Algorithm 5.
Input: A known chordal skeleton $G$ which is acyclically orientable with indegree $d$, a rooted clique tree $\mathcal{T}_{C_{r}}$ of $G$, hyperparameter $\eta$, a list of samples $[x^{(1)},\ldots,x^{(t)}]$ from $P^{*}$, time step $t$.
Output: $3$-dimensional table $\mathsf{Table}$.
1
[…]
4for every leaf node $C\in V(\mathcal{T}_{C})$ do
5       for every $\mathcal{O}_{C}\in\mathsf{AO}_{d}(C)$ do
6             for every $v\in V(G_{C})$ do
7                   Compute $\texttt{wt}_{C}(v,\mathcal{O}_{C},t)$ from the samples according to Equation 10.
8            $\mathsf{Table}[C,\mathcal{O}_{C},t]=\prod_{v\in C}\texttt{wt}_{C}(v,\mathcal{O}_{C},t)$ according to Equation 11.
9// The root is at level 00, and the lowest leaf is at level $\mathsf{NumberOfLevels}$.
10 for $j=\mathsf{NumberOfLevels}-1$ to 00 do
11       $\mathcal{S}\leftarrow$ the nodes at level $j$ in $\mathcal{T}^{G}$.
12       for every non-leaf $C\in\mathcal{S}$ do
13             Let the children of $C$ be $C_{1},\ldots,C_{\ell}$, which are at level $j+1$ by definition.
14             for every $\mathcal{O}_{C}\in\mathsf{AO}_{d}(\mathsf{Link}(C))$ do
15                   Compute $\mathsf{Table}[C,\mathcal{O}_{C},t]$ according to 5.11
16Return $\mathsf{Table}$.
Algorithm 5 $\mathsf{CountChordalDist}$ ($G,\mathcal{T}_{C_{r}},\eta,\mathsf{SampleList}=[x^{(1)},\ldots,x^{(t)}],t$)
Input: A chordal skeleton $G$ which is acyclically orientable with indegree $d$, hyperparameter $\eta>0$, a list of samples $x^{(1)},\ldots,x^{(t)}$ from distribution $P^{*}$ on $[k]^{n}$, time step $t\in[T]$
Output: A DAG with skeleton $G$.
1
2
3Construct a clique tree $\mathcal{T}^{G}$ of $G$ using the algorithm of [65].
[…]
7Call the counting algorithm $\mathsf{CountChordalDist}$ ($G,\mathcal{T}^{G},\eta,\mathsf{SampleList}$) to obtain a 3-d DP table $\mathsf{Table}$.
8Push $(r,\emptyset)$ to the $\mathsf{Stack}$
9while Stack is not empty do
10       Pop $(B,\mathcal{O}_{p})$ from $\mathsf{Stack}$, where $B$ is a node in $\mathcal{T}^{G}$ and $\mathcal{O}_{p}\in\mathsf{AO}_{d}(\mathsf{Link}(\mathsf{Pa}_{\mathcal{T}^{G}}(B)))$ if $B\neq r$
11      for every orientation $\mathcal{O}\in\mathsf{AO}(\mathsf{Link}(B),\mathcal{O}_{p})$ do
12
$P_{B}(\mathcal{O})\leftarrow\frac{\mathsf{Table}[B,\mathcal{O},t]}{\sum\limits_{\mathcal{O}^{\prime}\in\mathsf{AO}(\mathsf{Link}(B),\mathcal{O}_{p})}\mathsf{Table}[B,\mathcal{O}^{\prime},t]}$
13      Sample $\mathcal{O}_{B}\in\mathsf{AO}(\mathsf{Link}(B))$ from the distribution $P_{B}$
14      Fix the orientation of $\mathsf{Link}(B)$ to be $\mathcal{O}_{B}$
15      for each child $C$ of $B$ in $\mathcal{T}^{G}$ do
16             Push $(C,\mathcal{O}_{B})$ to $\mathsf{Stack}$
17Return all the orientations of the edges in $G$.
Algorithm 6 $\mathsf{SamplingChordalDist}$ $G,d,\eta,\mathsf{SampleList}=[x^{(1)},\ldots,x^{(t)}],t$)
Input :    Chordal skeleton $G$, indegree parameter $d$, $\mathcal{N}=\prod_{i\in[n]}\prod_{S\subseteq{\rm Nbr}_{G}(i)}\mathcal{N}^{X_{i}\mid X_{S}}_{\varepsilon/n}$, hyperparameter $\eta>0$, $T$.
Output :    A chordal-structured distribution $\widehat{P}$.
1
2Sample $\mbox{Sample }t\leftarrow\{1,\ldots,T\}\mbox{ uniformly at random.}$.
[…]
4for $s\leftarrow 1$ to $t$ do
5       $\mbox{Observe sample }x^{(s)}\sim P^{\ast}$.
6      $\mathsf{SampleList}\leftarrow\mathsf{SampleList}\cup\{x^{(s)}\}$.
7Call $\mathsf{SamplingChordalDist}$ ($G,d,\eta,\mathsf{SampleList}$,t) to obtain a DAG $D$.
8For each $i\in[n]$, sample $\widehat{p}_{i}$ as described in text.
9return the Bayes net $\widehat{P}=(\widehat{p}_{1},\dots,\widehat{p}_{n})$ on the DAG $D$.
Algorithm 7 $\mathsf{LearnChordalDist}$ (RWM-based proper learning of chordal-structured distributions)
Once $\mathsf{Table}[\cdot,\cdot,t]$ has been constructed, we can sample an orientation of the chordal graph $G$ at time step $t$. The idea is to go down from the root of $\mathcal{T}_{C_{r}}$ to its leaves, level-by-level, with each iteration orienting the link of a maximal clique while being consistent with the orientations sampled so far. For each clique $C\in V(\mathcal{T}_{C_{r}})$ with parent $C_{p}$, we suppose that the edges in $\mathsf{Link}(C_{p})$ have been oriented fully. We use the table $\mathsf{Table}[\cdot,\cdot,t]$ to sample in a consistent way with the already oriented edges. To argue correctness, we again appeal to 5.6 to see that once a parent clique has been oriented, its children can be oriented independently as long as they are consistent with the parent’s orientation. Algorithm 6 gives the pseudocode.
Suppose $D$ is the orientation of $G$ that is sampled by $\mathsf{SamplingChordalDist}$ for a uniformly chosen value of $t\in[T]$. Then, we can readily sample a Bayes net on $D$ by following the same strategy described in the proof of LABEL:{lem:ewa_poly_tree}. Namely, for each $i\in[n]$, we sample $\widehat{p}_{i}(y_{i},y_{\mathsf{Pa}_{D}(i)})$ from $\mathcal{N}_{\varepsilon/n}^{X_{i}\mid X_{\mathsf{Pa}_{D}(i)}}$ independently, where:
$\Pr[\widehat{p}_{i}=p]\propto\prod_{s=1}^{t}\exp_{\eta}\log p\left(x^{(s)}_{i},x^{(s)}_{\mathsf{Pa}_{D}(i)}\right).$
Finally, the sampled Bayes net $\widehat{P}$ is obtained by letting the conditional probability distribution at each node $i$ be $\widehat{p}_{i}$. The proper learning algorithm is summarized as pseudocode in Algorithm 7. The improper learning algorithm is similar, but it uses EWA in place of RWM, and so, it samples a random $\widehat{P}$ when generating samples instead of during the learning phase.
Acknowledgements
AB and PGJ’s research were supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. AB and SS were supported by National Research Foundation Singapore under its NRF Fellowship Programme (NRF-NRFFAI1-2019-0002). AB was additionally supported by an Amazon Research Award and a Google South/Southeast Asia Research Award. SG’s work is partially supported by the SERB CRG Award CRG/2022/007985. NVV’s work was supported in part by NSF CCF grants 2130608 and 2342244 and a UNL Grand Challenges Catalyst Competition Grant.
We would like to thank Debojyoti Dey, a Ph.D. student at IIT Kanpur, for discussions regarding robust learning algorithms in high dimensions. AB would also like to thank Daniel Beaglehole for a short meeting which seeded the idea for this work.
Appendix A Lower bound for learning tree-structured distributions
In this section, we show a lower bound of $\Omega(\frac{n}{\varepsilon})$ holds for our tree structure learning results. This follows from [8]. We are including it here for completeness.
Theorem A.1.
Given sample access to an unknown tree-structured discrete distribution defined over $[k]^{n}$, and $\varepsilon>0$ be a parameter. With probability at least $9/10$, $\Omega(\frac{n}{\varepsilon})$ samples are necessary for any realizable PAC learner for learning $\mathcal{C}^{\textsc{Tree}}$.
We will prove the above result for $k=2$, that is, for distributions defined over $\{0,1\}^{n}$. The lower bound immediately extends to distributions over $[k]^{n}$.
We will first prove the above result for $n=3$ nodes. Then we will extend the result for $n=3\ell$, for some positive integer $\ell$.
Definition A.2 (Hellinger distance).
Let $P$ and $Q$ be two probability distributions defined over the same sample space $\mathcal{D}$. The Hellinger distance between $P$ and $Q$ is defined as follows:
$\mathsf{H}(P,Q)=\sqrt{\frac{1}{2}\sum_{x\in\mathcal{D}}\left(\sqrt{P(x)-\sqrt{Q(x)}}\right)^{2}}$
Fact A.3.
Let $P$ and $Q$ be two probability distributions defined over the same sample space $D$. Then the following holds:
$\mathsf{H}(P,Q)\leq\sqrt{\frac{1}{2}\mathsf{D_{KL}}(P||Q)}$
We will also need the following folklore result.
Fact A.4.
Let $P$ be an unknown discrete distribution defined over $\{0,1\}^{3}$, and $\varepsilon>0$ be a parameter. In order to output a distribution $\widehat{P}$ defined over $\{0,1\}^{3}$ such that $\mathsf{H}(P,\widehat{P})\leq\sqrt{\varepsilon}$, $\Omega(\frac{1}{\varepsilon})$ samples from $P$ are necessary.
Lemma A.5.
There exist three tree-structured distributions $P_{1},P_{2},P_{3}$ defined over $\{0,1\}^{3}$, such that any algorithm that that can learn any distribution $P_{i}$ in up to $\varepsilon$-KL-distance, that is, can output a distribution $\widehat{P}_{i}$ such that $\mathsf{D_{KL}}(P_{i}||\widehat{P}_{i})\leq\varepsilon$, requires $\Omega(\frac{1}{\varepsilon})$ samples from $P_{i}$, for any $i\in[3]$.
$P_{1}$ has three nodes $X_{1},Y_{1},Z_{1}$, $P_{2}$ has three nodes $X_{2},Y_{2},Z_{2}$ and $P_{3}$ has three nodes $X_{3},Y_{3},Z_{3}$. In $P_{1}$, $(Y_{1},Z_{1})$ takes values uniformly at random between $(0,0)$ and $(1,1)$ and $X_{1}$ copies $Y_{1}$ and $Z_{1}$ with probability $(1-\varepsilon)$, and with the remaining probability, takes a value from $\mathsf{Ber}(\frac{1}{2})$. $P_{2}$ and $P_{3}$ are defined in a similar fashion where we set $X_{2}=Z_{2}$ and $X_{3}=Y_{3}$, respectively.
We will use the following results from [8].
Claim A.6 (See Fact 7.4 (i) and Lemma 7.5 (i) of [8]).
  • (i)
    $\mathsf{H}(P_{i},P_{j})\geq 10\sqrt{\varepsilon}$ for any $i\neq j\in[3]$.
  • (ii)
    Every tree $\mathcal{T}$ on $3$ vertices is not $\Theta(\varepsilon\log\frac{1}{\varepsilon})$-approximate for any of $P_{1},P_{2}$ and $P_{3}$.
Proof of A.5.
We will prove this by contradiction. Suppose there exists an algorithm $\mathcal{A}$ that can learn any distribution up to $\varepsilon$-KL-distance using $o(\frac{1}{\varepsilon})$ samples, by outputting the pmf of $\widehat{P}$. In that case, we will argue that using $\mathcal{A}$, we can distinguish between $P_{1},P_{2}$ and $P_{3}$.
Let us denote the unknown distribution be $P$, and suppose given sample access to $P$, $\mathcal{A}$ takes $o(1/\varepsilon)$ samples from $P$ and outputs a distribution $\widehat{P}$ such that $\mathsf{H}(P,\widehat{P})\leq\sqrt{\varepsilon}$. This implies that given sample access to either of $P_{i}$ with $i\in[3]$, $\mathcal{A}$ can output such a $\widehat{P}$ such that $\mathsf{H}(P_{i},\widehat{P})\leq\sqrt{\varepsilon}$, which contradicts A.4. Using A.3, we are done with the proof of the lemma. ∎
Now we are ready to prove A.1.
Proof of A.1.
Let us assume that $n=3\ell$ for some positive integer $\ell$. We will divide the $n$ variables into $\ell$ blocks, each of size $3$ nodes. We will prove this by contradiction. We will define a distribution $P$ over $\{0,1\}^{n}$, where the $i$-th block is chosen uniformly to be either $P_{1}$ or $P_{2}$.
Suppose there exists an algorithm that can take $o(\frac{n}{\varepsilon})$ samples and outputs an $\varepsilon$-approximate tree $T$ of $P$, with probability at least 9/10. Since we have chosen each block independently, $T$ is a disjoint union of $T_{1},\ldots,T_{\ell}$. Using A.6, and setting $\varepsilon$ as $\varepsilon/\ell$, we can say that each $T_{i}$ is not $\Theta(\varepsilon/\ell)$-approximate with probability at least $2/3$. Thus, using Chernoff bound, we can say that at least $\frac{\ell}{100}$ trees are not $\frac{C\varepsilon}{\ell}$ approximate. Thus, $\mathsf{D_{KL}}(P||\widehat{P})\geq\varepsilon$ for a suitable constant $C$ for any tree-structured distribution $\widehat{P}$. This completes the proof of the theorem. ∎
Appendix B Learning Bayesian networks with bounded vertex cover
In this section, we show that bayesian networks can be learned efficiently if the size of the vertex cover of the associated moralized graph is bounded. Before proceeding to present our result, we need some notations.
Let $G=(V,E)$ be DAG with the vertex set $V$ and the edge set $E$. Here each vertex $v\in V$ corresponds to a variable and the edges in $E$ encode the conditional independence relations between the nodes of $V$. Let $\mathcal{D}$ denote the set of all possible DAG on the vertex set $V$. Moreover, for every DAG $G$, let us associate it with a non-negative weight function $f:G\rightarrow\mathbb{R}^{+}\cup\{0\}$ which encodes how well $G$ fits a given dataset. Now we are ready to define the notion of modular weight function, which will be crucially used in our proofs.
Definition B.1 (Modular weight function).
Let $G(V,E)$ be a DAG with vertex set $V$ and edge set $E$. Moreover, let $f:G\rightarrow\mathbb{R}^{+}\cup\{0\}$ denotes the weight function associated with $G$. The weight function $f$ is said to be a modular weight function if it has the following form:
$f(G)=\prod_{v\in V}f_{v}(G_{v})$ (14)
Now we define the notion of the vertex cover number of a DAG below.
Definition B.2 (Vertex cover number of a DAG).
Given a DAG $G$, let $G_{M}$ denote the undirected moralized graph corresponding to $G$. The vertex cover number $\tau(G)$ of $G$ is the size of the smallest vertex cover of $G_{M}$. Moreover, for some integer $\ell$, the set of DAGs $G$ such that $\tau(G)\leq\ell$ is denoted as $\mathcal{DAG}^{(\ell)}$.
Now we are ready to state the main result that we will be proving in this section.
Theorem B.3.
Let $P^{*}$ be an unknown discrete distribution on $[k]^{n}$ defined over a DAG $G$ such that the size of the vertex cover of the moralized graph $G_{M}$ corresponding to $G$ is bounded by an integer $\ell$. Moreover, let $\mathcal{DAG}^{(\ell)}$ be the family of distributions over $[k]^{n}$ that can be defined as Bayes nets over DAGs whose vertex cover size of the associated moralized graph is bounded by $\ell$. Given sample access from $P^{*}$, and a parameter $\varepsilon>0$, there exist the following:
  • (i)
    An agnostic PAC-learner for $\mathcal{DAG}^{(\ell)}$ using $O(n^{4}k^{2\ell+2}\varepsilon^{-4}\log(1/\delta))$ samples and $O(\exp(\ell)\mathrm{poly}(n))$ running time that is improper and returns an efficiently samplable mixture of distributions from $\mathcal{DAG}^{(\ell)}$.
  • (ii)
    An agnostic PAC-learner for $\mathcal{DAG}^{(\ell)}$ using $O(n^{3}k^{\ell+1}\delta^{-2}\varepsilon^{-2}\log(nk/\varepsilon))$ samples and runs in time $O(\exp(\ell)\mathrm{poly}(n))$ that is proper and returns a distribution from $\mathcal{DAG}^{(\ell)}$.
In order to prove the above theorem, we will be using the following result from [47] which states that $\mathcal{DAG}^{(\ell)}$ can be sampled efficiently.
Theorem B.4 (Theorem $13$ of [47] restated).
There exists a randomized algorithm that can sample from $\mathcal{DAG}^{(\ell)}$ with weight proportional to $f(G)$ efficiently in expected sampling time $4^{\ell}n^{O(1)}$.
Note that the above result gives a polynomial time sampling algorithm for $\mathcal{DAG}^{(\ell)}$ when $\ell$ is bounded by a constant.
Interestingly, our loss function is also a modular weight function.
Claim B.5.
Our loss function $\exp_{\beta}\log\left(\prod_{s=1}^{t}P(x^{(i)})\right)$ is a modular weight function.
This follows from the fact that the distribution $P^{*}$ defined over the DAG $G$ factorizes: see Equation 1 in 2.3. Now we are ready to present the proof of our main result.
Proof of B.3.
  • (i)
    We will be using the EWA algorithm (Algorithm 1) for this purpose. In each round of the algorithm, we will be using the algorithm from B.4. The sample complexity (aka. number of rounds of EWA algorithm) follows from the guarantee of EWA algorithm for learning Bayesian networks (see 3.12). Since in each round, we will be calling the algorithm corresponding to B.4, the running time of our algorithm will be $O(\exp(\ell)\mathrm{poly}(n))$.
  • (ii)
    We will be calling RWM algorithm (Algorithm 2) here. Similar to the above, in each round, we will be using the algorithm from B.4. The sample complexity (aka. number of rounds of RWM algorithm) follows from the guarantee of RWM algorithm for learning Bayesian networks (see 3.13). Since in each round, we will be calling the algorithm corresponding to B.4, the running time of our algorithm will be $O(\exp(\ell)\mathrm{poly}(n))$.
Appendix C Learning Tree-structured distributions
Full proof of 4.3
As in Section 4, given a vertex set $V=[n]$, $\mathcal{G}_{\rm T}$ denotes the set of all spanning out-arborescences rooted at node $1$. $\mathcal{C}^{\rm\textsc{Tree}}$ denotes the set of all tree-structured distributions on $[k]^{n}$. $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$ denotes the finite set of distributions (see 4.2) that discretize $\mathcal{C}^{\rm\textsc{Tree}}$ with error $\varepsilon$ in $\mathsf{D_{KL}}$. Each tree-structured distribution $P$ in $\mathcal{N}^{\rm\textsc{Tree}}_{\varepsilon}$ can be factored as $(G,V_{P})$ where $G$ is the underlying rooted arborescence and $V_{P}=(p_{e}:e\in E(G),p_{1})$ is the vector of functions where $p_{e}(x_{i},x_{j})$ for $e=(i,j)$ corresponds to $\Pr_{X\sim P}(X_{j}=x_{j}\mid X_{i}=x_{i})$ and $p_{1}(x_{1})$ corresponds to $\Pr_{X\sim P}(X_{1}=x_{1})$. Using the notation of 3.6, $p_{1}\in\mathcal{N}^{X_{1}}_{\varepsilon/n}$ and for $e=(i,j)$, $p_{e}\in\mathcal{N}^{X_{j}\mid X_{i}}_{\varepsilon/n}$. For simplicity, we will call these sets $\mathcal{N}^{1}$ and $\mathcal{N}^{e}$ respectively in this section. Node that $|\mathcal{N}^{1}|,|\mathcal{N}^{e}|\leq\left(\frac{nk}{\varepsilon}\right)^{k^{2}}$, which is polynomial in $n$ and $1/\varepsilon$ for constant $k$.
In the proof of 4.3, we derive the following expression for $\omega_{t}(P)$, the weight assigned to a tree-structured distribution $P\cong(G,(p_{e}:e\in E(G),p_{1}))$ in the discretized set $\mathcal{N}$ by the EWA/RWM algorithms at step $t$.
$\omega_{t}(P)$ $=\frac{\prod_{s<t}\exp_{\eta}\log p_{1}(x^{(s)}_{1})}{\sum_{q_{1}\in\mathcal{N}^{1}}\prod_{s<t}\exp_{\eta}\log q_{1}(x^{(s)}_{1})}\cdot\frac{\prod_{e\in E(G)}\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}{\sum_{H\in\mathcal{G}_{\rm T}}\prod_{e\in E(H)}\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}\cdot$
$\qquad\qquad\prod_{e\in E(G)}\frac{\prod_{s<t}\exp_{\eta}\log p_{e}(x^{(s)}_{e})}{\sum_{q_{e}\in\mathcal{N}^{e}}\prod_{s<t}\exp_{\eta}\log q_{e}(x^{(s)}_{e})}.$ (15)
This gives the following high-level algorithm for sampling from the EWA/RWM distribution.
  • (i)
    Sample $p_{1}$ from the distribution over $\mathcal{N}^{1}$ with the following probability:
    $\Pr(p_{1})\triangleq\frac{\left(\prod_{s<t}p_{1}(x^{(s)}_{1})\right)^{\eta}}{\sum_{q_{1}\in\mathcal{N}^{1}}\left(\prod_{s<t}q_{1}(x^{(s)}_{1})\right)^{\eta}}.$
  • (ii)
    Sample a spanning arboresence $G\in\mathcal{G}_{\rm T}$ (rooted at node $1$), where $\Pr(G)\propto\prod\limits_{e\in E(G)}\mathbf{w}(e)$ and $\mathbf{w}(e)\triangleq\sum\limits_{q_{e}\in\mathcal{N}^{e}}\left(\prod_{s<t}q_{e}(x^{(s)}_{e})\right)^{\eta}$ for any $e=(i,j)\in[n]^{2}$. This can be done in polynomial time (even though there are exponentially many such arborescences) by applying Tutte’s matrix-tree theorem to weighted digraphs [10, 16, 75, 37]. We describe this approach in the algorithm SamplingArborescence, which we describe and analyze later. We invoke
    $G\leftarrow\textsc{SamplingArborescence}(G_{0}\leftarrow K_{n},w\leftarrow\mathbf{w},r\leftarrow 1),$
    where $K_{n}$ is the complete graph on the vertex set $[n]$ viewed as a digraph, which returns $G\in\mathcal{G}_{\rm T}$ (rooted at $1$) with the required sampling probability.
  • (iii)
    For each $e\in E(G)$ (fixed in the preceding step), $p_{e}$ is sampled from a product distribution over $\mathcal{N}^{e}$, with the following probability:
    $\Pr(p_{e})\triangleq\frac{\left(\prod_{s<t}p_{e}(x^{(s)}_{e})\right)^{\eta}}{\sum_{q_{e}\in\mathcal{N}^{e}}\left(\prod_{s<t}q_{e}(x^{(s)}_{e})\right)^{\eta}}.$
Once $P$ is sampled, a sample from $P$ can be generated in polynomial time as $P$ is tree-structured.
Correctness
The correctness of the above algorithm follows from considering each factor of Appendix C and noting that the steps (i)-(iii) described above are independent.
Running Time
In Step (i), using memoization between time steps appropriately, $p_{1}$ can be sampled in time $O(|\mathcal{N}^{1}|)\leq(nk/\varepsilon)^{O(k)}$. The idea is to keep $\mathsf{DPT}[p]=\left(\prod_{s<t-1}p(x^{(s)}_{1})\right)^{\eta}$ for each $p\in\mathcal{N}^{1}$ from the sampling procedure at step $(t-1)$, and multiply $\mathsf{DPT}[p]$ with $p(x_{t-1,1})^{\eta}$ before step $t$ to get $\left(\prod_{s<t}p(x^{(s)}_{1})\right)^{\eta}$. A similar idea can be used in step (iii) to sample each $p_{e}$ ($n-1$ of them) in time $O(|\mathcal{N}^{e}|)\leq(nk/\varepsilon)^{O(k^{2})}$. The SamplingArborescence call in this case (invoked with $K_{n}$) requires $O(n^{5})$ time, which gives an overall time complexity which is $\mathsf{poly}(n,1/\varepsilon)$ for constant alphabet size $k$.
The SamplingArborescence algorithm
Preliminaries
Let $G=([n],E,w)$ be a connected weighted directed graph on $n$ vertices $\{1,\ldots,n\}$ with $m$ edges $e_{1},\ldots,e_{m}$. $w:E\rightarrow\mathbb{R}_{>0}$ is a positive weight function. Let $A_{w}(G)$ denote the $n\times n$ weighted vertex adjacency matrix of $G$. For every $i,j\in[n]$, $A_{w}(G)$ is defined as follows:
$[A_{w}(G)]_{i,j}=\begin{cases}w(e_{k}),&\mbox{if $e_{k}$ is the directed edge from $i$ to $j$}\\ 0,&\mbox{if there is no directed edge from $i$ to $j$}\end{cases}$
Let $N_{\mathrm{in},w}(G)$ be the weighted inward incidence matrix of order $m\times n$ defined as follows: For every $i\in[n]$ and $k\in[m]$, we have:
$[N_{\mathrm{in},w}(G)]_{k,i}=\begin{cases}\sqrt{w(e_{k})},&\text{if directed edge $e_{k}$ points to vertex $i$}\\ 0,&\text{otherwise}\end{cases}$
Similarly, we can define the weighted outward incidence matrix $M_{\mathrm{out},w}(G)$ of order $n\times m$ as follows, for $i\in[n]$ and $k\in[m]$:
$[M_{\mathrm{out},w}(G)]_{i,k}=\begin{cases}\sqrt{w(e_{k})},&\text{if directed edge $e_{k}$ points from vertex $i$}\\ 0,&\text{otherwise}\end{cases}$
The indegree matrix $D_{\mathrm{in},w}(G)$ is a $n\times n$ diagonal matrix such that, for all $i\in[n]$, $[D_{\mathrm{in},w}(G)]_{i,i}$ is equal to the sum of the weights of all incoming edges to vertex $i$. Similarly, the outdegree matrix $D_{\mathrm{out},w}(G)$ is a $n\times n$ diagonal matrix such that for all $i\in[n]$, $[D_{\mathrm{out},w}(G)]_{i,i}$ is equal to the sum of the weights of all outgoing edges from vertex $i$.
Now we can define the Laplacian matrices $L_{1,w}(G)$ and $L_{2,w}(G)$ associated with digraph $G$ as follows:
$L_{1,w}(G)\triangleq D_{\mathrm{in},w}(G)-A_{w}(G),\ \mbox{and}\ L_{2,w}(G)=D_{\mathrm{out},w}(G)-A^{\top}_{w}(G).$ (16)
For all these associated matrices, we omit $G$ and denote the matrix $A_{w}(G)$ as $A_{w}$ etc. when $G$ is clear from the context. Now we have the following relationships between the Laplacian and incidence matrices:
Claim C.1 (Equation ($10$) of [37]).
  • (i)
    $L_{1,w}=(N^{\top}_{\mathrm{in},w}-M_{\mathrm{out},w})N_{\mathrm{in},w}$.
  • (ii)
    $L_{2,w}=(M_{\mathrm{out},w}-N^{\top}_{\mathrm{in},w})M^{\top}_{\mathrm{out},w}$.
For a vertex $r\in[n]$, let $L^{r}_{1,w}$ be the $(n-1)\times(n-1)$ matrix obtained by removing the $r$-th row and $r$-th column of $L_{1,w}$. Similarly, we also define $N^{r}_{\mathrm{in},w}$ and $M^{r}_{\mathrm{out},w}$. Then we have the following:
Claim C.2 (Equation ($11$) of [37]).
  • (i)
    $L^{r}_{1,w}=((N^{r}_{\mathrm{in},w})^{\top}-M^{r}_{\mathrm{out},w})N^{r}_{\mathrm{in},w}$.
  • (ii)
    $L^{r}_{2,w}=((M^{r}_{\mathrm{out},w}-(N_{\mathrm{in},w}^{r})^{\top}))(M_{\mathrm{out},w}^{r})^{\top}$.
Definition C.3 (Weight of a graph).
Let $G=(V,E,w)$ be a weighted (directed or undirected) graph, where $w:E\rightarrow\mathbb{R}_{>0}$. Let $H$ be any subgraph of $G$. The weight of graph $H$ is defined as the product of the weights of its edges, i.e.
$w(H)=\prod_{e\in E(H)}w(e).$
Definition C.4 (Contraction and deletion).
If $G=(V,E,w)$ is a weighted undirected graph (not necessarily simple) and $e=\{i,j\}\in E$, the graph $G/e$ obtained by contracting edge $e$ is a weighted graph on $V\setminus\{i,j\}\cup\{\langle i,j\rangle\}$ where
  • Vertices $i$ and $j$ are removed, and a new vertex $\langle i,j\rangle$ is added.
  • All edges $\{i,j\}$ (in case of parallel edges) are removed in $G/e$.
  • Edges $\{u,v\}$ where $u,v\in V\setminus\{i,j\}$ are preserved in $G/e$ with the same weight.
  • Every edge $\{u,v\}\in E$ where $u\in\{i,j\}$ and $v\in V\setminus\{i,j\}$ becomes an edge $\{\langle i,j\rangle,v\}$ in $G/e$ with the same weight.
Similarly the graph $G\setminus e$ obtained by deleting edge $e$ is just $(G\setminus e)=(V,E\setminus e,w\mid_{E\setminus e})$ where the edge $e$ is deleted (including all parallel edges) from $G$.
Definition C.5 (Contraction edge mapping).
With the above definition of contraction (with parallel edges kept), if $G$ is a graph and $G^{\prime}=G\setminus e$ for $e=\{i,j\}\in E(G)$, we can map each edge $e\in E(G^{\prime})$ injectively to an edge $e\in E(G)$ that caused its inclusion in $G^{\prime}$. We denote this mapping by $f_{G^{\prime}:G}:E(G^{\prime})\rightarrow E(G)$.
The following theorem is a generalization of Tutte’s matrix-tree theorem to weighted graphs.
Lemma C.6 (Theorem 3 of [37]).
Let $G=(V,E,w)$ be a weighted directed graph on $n$ vertices $[n]$ with positive weight function $w:E\rightarrow\mathbb{R}_{>0}$. Then, for any $r\in[n]$, the sum of the weights of the weighted outgoing (incoming) spanning arborescences rooted at vertex $r$ is equal to $\det(L^{r}_{1,w})$ ($\det(L^{r}_{2,w})$, respectively).
Now the next corollary follows as any undirected graph can be viewed as a directed graph with edge $\{i,j\}$ corresponding to a pair of directed edges $(i,j)$ and $(j,i)$.
Corollary C.7.
If $G=K_{n}$ (the complete graph on $n$ vertices) and $w:[n]^{2}\rightarrow\mathbb{R}_{>0}$ is any positive weight function,
$\det(L^{1}_{1,w})=\sum_{G\in\mathcal{G}_{\rm T}}\prod_{e\in E(G)}w(e),$
where $\mathcal{G}_{\rm T}$, as defined earlier, is the set of all spanning out-arborescences on $[n]$ rooted at vertex $1$.
Input: $G_{0}$ a directed graph, $w:E(G_{0})\rightarrow\mathbb{R}_{\geq 0}$ weight function, $r\in V(G_{0})$ root node.
Output: Arborescence $G_{T}\in\mathcal{A}_{G_{0},r}$ with $\Pr(\text{output }G_{T})\propto\prod_{e\in E(G_{T})}w(e)$.
1 Let $G_{T}\leftarrow\emptyset$.
2 Choose an arbitrary ordering of the $m$ edges of $G_{0}$, and let $\mathsf{EdgeList}\leftarrow[e_{1},\ldots,e_{m}]$.
3 Let $\mathsf{ContractionMapping}[e_{i}]\leftarrow e_{i}$ for each $i\in[m]$.
4 Let $j\leftarrow 0$.
5 while $r$ has an outgoing edge in $G_{j}$ do
6       Let $e=(r,x)$ be the first remaining outgoing edge.
7       Compute $p_{e}\leftarrow\frac{\det(L^{r}_{1,w}(G_{j}\setminus e))}{\det(L^{r}_{1,w}(G_{j}))}$.
8       Sample $u\sim\mathsf{Unif}((0,1])$.
9       if $u\leq p_{e}$ (with probability $p_{e}$) then
10             $G_{j+1}\leftarrow G_{j}\setminus e$ (delete edge $e$).
11             $j\leftarrow j+1$.
12
13      else
14             $G_{j+1}\leftarrow G_{j}/e$ (contract edge $e$).
15             Add $\mathsf{ContractionMapping}[e]$ to $G_{T}$ (this will be an edge of $G_{0}$).
             /* Reconstruct ContractionMapping for mapping $E(G_{j+1})$ to $E(G_{0})$. */
16             $\mathsf{CM}^{\prime}\leftarrow\textrm{empty dictionary}$.
17             for each edge $e\in E(G_{j+1})$ do
18                   $\mathsf{CM}^{\prime}[e]\leftarrow\mathsf{ContractionMapping}[f_{G_{j+1}:G_{j}}(e)]$ (see C.5).
19
20            $\mathsf{ContractionMapping}\leftarrow\mathsf{CM}^{\prime}$
21             $r\leftarrow\langle r,x\rangle\in V(G_{j+1})$ (the newly-contracted vertex).
22             $j\leftarrow j+1$.
23
return $G_{T}$.
Algorithm 8 The SamplingArborescence algorithm
The algorithm
If $G$ is a graph and $v\in V(G)$ be a vertex of $G$, let $\mathcal{A}_{G,v}$ denote the set of out-arborescences on $V(G)$ which are subgraphs of $G$ (viewed as a digraph), rooted at $v$, and span the set of nodes reachable from $v$. Note that this set will not be empty unless $v$ is an isolated vertex.
Definition C.8 (Product arborescence distribution).
Let $G=(V,E,w)$ be a weighted graph where $w:E\rightarrow\mathbb{R}_{\geq 0}$ is a weight function on the edges. For any $v\in V(G)$, a distribution $\mu$ on $\mathcal{A}_{G,v}$ is said to be the product arborescence distribution if
$\Pr_{G\sim\mu}(G)\propto\prod_{e\in G}w(e)\mbox{ for all $G\in\mathcal{A}_{G,v}$}.$
Claim C.9.
Suppose $G=(V,E,w)$ is a weighted digraph with non-negative weights, $r\in V(G)$ and $\mu$ is a product arborescence distribution on $\mathcal{A}_{G,r}$. Then
$\Pr_{T\sim\mu}(e\in T)=1-\frac{\det(L^{r}_{1,w}(G\setminus e))}{\det(L^{r}_{1,w}(G))}.$
Proof.
Since $\Pr_{\mu}(T)\propto\prod_{e\in T}w(e)$, we have
$\Pr_{T\sim\mu}(e\in T)=\frac{\sum_{T\in\mathcal{A}_{G,r}\,:\,T\ni e}w(T)}{\sum_{T\in\mathcal{A}_{G,r}}w(T)}=1-\frac{\sum_{T\in\mathcal{A}_{G,r}\,:\,T\not\ni e}w(T)}{\sum_{T\in\mathcal{A}_{G,r}}w(T)}=1-\frac{\det(L^{r}_{1,w}(G\setminus e))}{\det(L^{r}_{1,w}(G))},$
where the first equality follows from C.8 and C.3, and the third equality follows from applying C.6 to $G\setminus e$ and $G$. ∎
Now we are ready to prove the correctness of the algorithm SamplingArborescence.
Claim C.10.
There is an algorithm SamplingArborescence that, if invoked with input $(G,w,r)$ for a digraph $G$, $w:E(G)\rightarrow\mathbb{R}_{\geq 0}$, and $r\in V(G)$, returns a random arborescence $G_{T}$ sampled from the product arborescence distribution on $\mathcal{A}_{G,r}$. The algorithm runs in time $O(|E||V|^{3})$ in the real RAM model.
Proof.
The pseudocode for the algorithm is given as Algorithm 8. Observe that, if $G_{T}\in\mathcal{A}_{G,r}$, then it is actually sampled from the product arborescence distribution (C.8) by C.9, since each edge $e$ is added to $G_{T}$ with the correct probability $1-p_{e}$.
It remains to ensure that $G_{T}$ is indeed a spanning arborescence. Let $r_{j}$ denote the value of $r$ used in iteration $j$ of the while loop (lines 5-20). $r_{j}\in V(G_{j})$ by construction. Let $\mathcal{S}(r_{j})\subseteq V(G_{0})$ denote $\{r_{j}\}$ if $V(G_{j})=V(G_{0})$ (no contractions) and the set of nodes in $G_{0}$ corresponding to the contracted-vertex in $G_{j}$ otherwise.
By construction, when $|\mathcal{S}(r_{j})|>1$, the algorithm maintains (in $G_{T}$) an out-arborescence rooted at $r=r_{0}$ that spans $\mathcal{S}(r_{j})$. This is because when the algorithm selects an edge $e=(r_{j},x)\in E(G_{j})$ and takes $G_{j+1}=G_{j}/(r_{j},x)$, this edge will correspond to an edge $(v,x)$ in the original graph $G_{0}$ where $v\in\mathcal{S}(r_{j})$ and $x\not\in\mathcal{S}(r_{j})$. This will ensure that there are no (undirected) cycles formed when adding $(v,x)$ to $G_{T}$ and the rooted arborescence invariant will be maintained in iteration $j+1$ with $\mathcal{S}(r_{j+1})=\mathcal{S}(r_{j})\sqcup\{x\}$. Note that when an edge $e$ is selected for contraction in iteration $j$, the algorithm maintains a mapping $\mathsf{ContractionMapping}$ from the edges in $G_{j}$ to the edges in $G_{0}$ and adds the original edge (in $G_{0}$) to $G_{T}$. The other property of the algorithm is that it does not disconnect the graph — an edge $e$ which is a cut edge will be present in all arborescences and hence will have $p_{e}=0$, ensuring that it will never be selected for deletion (in line 9). So, the algorithm will terminate when $\mathcal{S}(r_{j})$ is the set of nodes reachable from $r=r_{0}$, and $G_{T}$ will have the required spanning arborescence.
For the time complexity bound, note that the while loop (lines 5-20) will run at most $|E|$ times, while the expensive step inside the while loop will be the two determinant computations in line 7 (to compute $p_{e}$). These can be done in $O(|V|^{3})$ time in the real RAM model, by Gaussian elimination etc. Contraction, deletion etc. require only $O(|V|+|E|)\leq O(|V|^{2})$ time. This gives us the $O(|E||V|^{3})$ bound. ∎